### v1

# Non-Linear Model-Based Sequential Decision-Making in Agriculture

Abstract
Sequential decision-making is central to sustainable agricultural management and precision agriculture, where resource inputs must be optimized under uncertainty and over time. However, such decisions must often be made with limited observations, whereas classical bandit and reinforcement learning approaches typically rely on either linear or black-box reward models that may misrepresent domain knowledge or require large amounts of data. We propose a family of nonlinear, model-based bandit algorithms that embed domain-specific response curves directly into the exploration-exploitation loop. By coupling (i) principled uncertainty quantification with (ii) closed-form or rapidly computable profit optima, these algorithms achieve sublinear regret and near-optimal sample complexity while preserving interpretability. Theoretical analysis establishes regret and sample complexity bounds, and extensive simulations emulating real-world fertilizer-rate decisions show consistent improvements over both linear and nonparametric baselines (such as linear UCB and $k$-NN UCB) in the low-sample regime, under both well-specified and shape-compatible misspecified models. Because our approach leverages mechanistic insight rather than large data volumes, it is especially suited to resource-constrained settings, supporting sustainable, inclusive, and transparent sequential decision-making across agriculture, environmental management, and allied applications. This methodology directly contributes to SDG 2 (Zero Hunger) and SDG 12 (Responsible Consumption and Production) by enabling data-driven, less wasteful agricultural practices.
Keywords: Sequential decision-making, agricultural statistics, nonlinear models, multi-armed bandits, resource optimization, regret
1 Introduction
Identifying the most effective crop management practice for a given site is a recurring challenge in agriculture. While field trials and expert recommendations provide useful starting points, the actual performance of management strategies such as different fertilizer rates, planting densities, or irrigation schedules can be highly variable due to differences in soil properties, previous management, and other biological or environmental factors (Hochman and Carberry,, 2011; Tilman et al.,, 2002). As a result, farmers and researchers must repeatedly decide which management actions to implement each season, often with limited prior information about their true effects. This problem exemplifies sequential decision-making under uncertainty, where the goal is to efficiently learn and optimize management choices over time.
In the statistical and machine learning literature, such problems are naturally framed as multi-armed bandit problems, dating back to Robbins, (1952). They have been widely applied in areas such as precision medicine (Lu et al.,, 2021), recommendation systems (Li et al.,, 2010), and, increasingly, precision agriculture (Gautron et al.,, 2022). The term originates from the analogy of a gambler choosing between several slot machines (“one-armed bandits”), each with unknown payout rates. At each round, the decision-maker selects an action/arm (e.g., a specific management practice) and observes its outcome (e.g., crop yield or profit). The central challenge is to balance exploration (trying new or uncertain options to learn about their effects) with exploitation (repeating choices that have performed well so far) in order to maximize cumulative reward or minimize loss over repeated decisions. Bandit algorithms provide principled, data-driven strategies to manage this tradeoff, making them well-suited to resource optimization tasks in agriculture, environmental management, and beyond. However, most existing bandit methods assume linear, generalized linear, or black-box (non-parametric) reward models that often fail to capture the mechanistic, nonlinear dose-response dynamics common in these domains.
In this work, we develop and analyze a new family of bandit algorithms that directly incorporate domain-specific nonlinear response models, offering greater interpretability and sample efficiency compared to classical linear or black-box approaches. Our approach is particularly valuable in settings where repeated decision-making is required but available data is limited and the underlying input-response relationship is nonlinear or complex.
Nonlinear models can encompass a wide range of mean functions. Although individual nonlinear models may be less flexible than general linear models (e.g., polynomials), they are often more parsimonious, interpretable, and tailored to domain-specific processes. Their interpretability stems from the ability to associate parameters with biologically meaningful phenomena. In agriculture, nonlinear models (Miguez et al.,, 2018) refer to mathematical equations that describe complex relationships between variables, such as crop yield response to fertilizer, temperature effects on plant growth, or pest spread–where the response is not linear but follows more intricate patterns (e.g., exponential or logistic). The adoption of such models is motivated not only by statistical fit but also by the need for transparent, explainable recommendations in sustainable resource management. Let us first provide a brief overview of research developments in sequential decision-making suitable for our discussion and motivation for developing new methodology and algorithms.
1.1 Background and related work
Multi-armed bandit algorithms (Berry and Fristedt,, 1985; Lattimore and Szepesvári,, 2018) provide a core statistical framework for sequential decision-making under uncertainty, where the key challenge is to balance exploration and exploitation. While the original bandit formulation assumed independent, identically distributed (i.i.d.) rewards, subsequent work has extended the framework to incorporate contextual information leading to the development of contextual bandit models (Slivkins,, 2014; Chu et al.,, 2011). In these settings, regression models such as linear (Abbasi-Yadkori et al.,, 2011), generalized linear models (GLMs) (Li et al.,, 2010; Filippi et al.,, 2010), and more flexible nonparametric approaches, including kernels, nearest neighbors, and neural networks (Yang and Zhu,, 2002; Rigollet and Zeevi,, 2010; Srinivas et al.,, 2010; Valko et al.,, 2013) enable data-driven, adaptive decision-making tailored to evolving arm/covariate information.
Classical nonlinear regression models, such as Mitscherlich (Dhanoa et al.,, 2022), Michaelis-Menten (Lopez et al.,, 2000), quadratic threshold models (Bélanger et al.,, 2000) and logistic response functions (Sepaskhah et al.,, 2011), are routinely used to capture crop yield response to fertilizer and other management variables. These models offer interpretable parameters tied to underlying biological processes, which is important for transparent and actionable recommendations (De Bruin and Pedersen,, 2008; Miguez et al.,, 2018).
Machine learning approaches, including random forests, neural networks, and regression trees, have become increasingly popular for crop yield prediction and resource optimization (Shahhosseini et al.,, 2019; Khaki et al.,, 2020; Khairunniza-Bejo et al.,, 2014; Jabed and Murad,, 2024). While powerful in large-data settings, these methods are typically supervised and static, lacking the sequential and adaptive capabilities needed for decision-making under uncertainty. Concerns also remain about their interpretability and suitability for low-data environments (Dobermann et al.,, 2022).
Within agriculture, adaptive experimentation and real-time learning strategies have emerged as promising approaches for site-specific management and resource efficiency (Gautron et al.,, 2022). Recent work has explored the use of bandit algorithms (Saikai et al.,, 2020, 2018; Huang,, 2025) for adaptive plot selection and management optimization, but most existing studies focus on linear models or generic black-box methods, seldom leveraging domain knowledge of nonlinear response behavior, especially valuable for smallholders and in data-scarce scenarios (Smith,, 2018).
It is well-known that model-based bandit and reinforcement learning approaches can be far more sample-efficient than model-free methods when the underlying reward mechanisms are known (Osband and Van Roy,, 2014; Sun et al.,, 2019). However, such approaches have received limited attention in agricultural decision-making from a precision-agriculture and sequential decision-making perspective, where established nonlinear models are often available but underutilized.
Our work addresses this gap by integrating domain-driven nonlinear models directly into the bandit framework. This enables adaptive, interpretable, and sample-efficient resource management, and advances the practical and theoretical toolkit for sustainable agriculture and related applications.
Our Contributions.
To address the limitations outlined above, this paper makes the following key contributions:
  • Nonlinear model-based bandit framework: We develop a general class of multi-armed bandit algorithms that explicitly incorporate domain-specific nonlinear response functions, such as Mitscherlich and logistic models, within the exploration-exploitation loop.
  • Theoretical guarantees: We provide regret and sample complexity results under both well-specified and shape-compatible misspecified models, and discussing the implications in the agricultural sequential decision-making domain.
  • Simulation studies emulating real-world practice: Through extensive simulations inspired by real-world fertilizer optimization under both well-specified and misspecified settings, we demonstrate the practical benefits of our approach over classical linear and nonparametric bandit methods, particularly in low-sample regimes.
  • Interpretability and domain relevance: By leveraging mechanistic models with interpretable parameters, our methods yield actionable recommendations for practitioners, supporting sustainable, transparent, and inclusive decision-making.
We begin by setting up the problem in Section 2. In Section 3, we lay out the proposed framework for nonlinear model-based bandits, in which we study three model-based algorithms: $\epsilon$-greedy, Upper Confidence Bound (UCB), and the ViOlin algorithm. Section 4 presents known results for parametric contextual bandits in both well-specified and misspecified settings, and derives sample complexity bounds for the specific nonlinear models considered by the ViOlin algorithm. We then conduct extensive simulation experiments emulating real-world yield production and fertilizer optimization scenarios, with the goal of profit maximization, especially in sample-limited settings. Finally, Section 6 concludes the paper.
2 Problem setup: Model-based nonlinear bandits
In agricultural systems, sequential decisions regarding inputs such as fertilizer rates, irrigation levels, or planting densities must be made under uncertainty, often with limited experimental data. These decisions involve selecting from a continuous (or discrete) set $\mathcal{X}$ of input levels (or arms) based on partially known relationships between inputs and outcomes, such as crop yield or profit. This setting naturally corresponds to a multi-armed bandit problem, where the goal is to sequentially identify input levels that maximize cumulative outcomes over time.
Formally, we consider a sequential decision-making scenario with a time horizon of $T$ rounds. In practice, each round can represent a growing season or field trial, with decisions made at the start and outcomes observed at harvest. At each round $t=1,\dots,T$, the decision-maker (agent) chooses an action (input level) $x_{t}\in\mathcal{X}$ from a predefined set of feasible treatments, such as nitrogen application rates. After choosing action $x_{t}$, the agent observes a noisy reward $y_{t}$, representing yield or profit, modeled as:
$y_{t}=f(x_{t};\beta^{*})+\epsilon_{t},$ (1)
where $f:\mathcal{X}\to\mathbb{R}$ is a known, structured nonlinear function parameterized by an unknown vector $\beta^{*}\in\mathbb{R}^{p}$, and $\epsilon_{t}$ is a mean-zero noise term. We assume $\epsilon_{t}$ is sub-Gaussian, i.e., there exists $\sigma>0$ such that for all $\lambda\in\mathbb{R}$,
$\mathbb{E}\left[e^{\lambda\epsilon_{t}}\right]\leq\exp\left(\frac{\lambda^{2}\sigma^{2}}{2}\right).$ (2)
This assumption ensures controlled tail behavior of the noise distribution, supporting reliable statistical concentration results.
The decision-maker seeks a policy/algorithm $\pi$ that selects actions at any given round based on historical observations to maximize cumulative expected reward. Equivalently, performance can be assessed through the cumulative regret, which quantifies the total expected loss incurred by not always selecting the optimal arm at each round. Specifically, the cumulative regret after $T$ rounds is defined as:
$R_{T}(\pi)=\sum_{t=1}^{T}\left\{f(x_{t}^{*};\beta^{*})-f(x_{t};\beta^{*})\right\},$ (3)
where $x_{t}^{*}=\arg\max_{x\in\mathcal{X}}f(x;\beta^{*})$ denotes the optimal treatment at round $t$, which is unknown to the decision-maker. Note that the cumulative regret $R_{T}(\pi)$ is a random quantity since it depends on the actions chosen by the policy $\pi$, which are themselves functions of the observed data.
The objective is to devise algorithms achieving sublinear regret, meaning $R_{T}(\pi)=o_{P}(T)$, so that the average regret per round diminishes over time. Such a property guarantees that the policy rapidly identifies near-optimal input levels, minimizing wasted resources and experimental costs. Other performance metrics more suitable for risk-averse decision-making have also been proposed, such as the Conditional Value-at-Risk (CVaR) of the regret, which emphasizes control of the worst-case outcomes in the tail of the regret distribution (Baudry et al.,, 2021; Wang et al.,, 2023).
To illustrate our methodology, we focus on fertilizer-yield relationships and specifically investigate four biologically motivated nonlinear models commonly employed in agricultural research:
  1. 1.
    Mitscherlich Model: $f(x;A,b,d)=d+A\left(1-e^{-bx}\right)$, capturing saturating yield response to fertilizer inputs.
  2. 2.
    Michaelis-Menten Model: $f(x;a,b,d)=d+\frac{ax}{b+x}$, describing saturating nutrient uptake or growth responses.
  3. 3.
    Quadratic Threshold Model:
    $f(x;a,b,c,x_{0})=\begin{cases}a+bx+cx^{2}&\text{if }x\leq x_{0}\\ a+bx_{0}+cx_{0}^{2}&\text{if }x>x_{0}\end{cases},$
    modeling symmetric or asymmetric yield responses that are often observed empirically, often with a threshold term to denote a plateau.
  4. 4.
    Logistic Dose-Response Model: $f(x;A,B,C,d)=d+\frac{A}{1+\exp\left(-B(x-C)\right)}$, frequently used to model threshold or inflection point responses to inputs.
It is important to emphasize here that these models offer interpretable parameters tied to agronomic phenomena (e.g., saturation, growth rate, efficiency), making them not only statistically efficient but also scientifically meaningful. For example, in the Mitscherlich model, where $A$ denotes the maximum additional yield achievable by the input, $b$ controls the rate of response or efficiency of the input, and $d$ is the baseline yield without the input. A larger value of $b$ indicates that the yield saturates more rapidly with increasing input, whereas a smaller $b$ corresponds to a more gradual response. These parameters are directly tied to agronomic processes, making the resulting policy recommendations both transparent and actionable for practitioners. Beyond these specific models, the proposed framework is broadly applicable to agricultural management problems. These include models for biomass accumulation over time, plant growth as a function of temperature, light-use efficiency, and nitrogen loss in response to rainfall, many of which are well captured by classical nonlinear models. A comprehensive list of such models and their applications is given by Miguez et al., (2018) and references therein. Detailed descriptions, parameter choices, and practical implementations of these models are presented explicitly in simulation studies (Section 5).
3 Algorithms for model-based nonlinear contextual bandits
Bandit algorithms address the fundamental challenge of sequential decision-making under uncertainty, where one must choose among several options (arms) over time to maximize cumulative reward. Recall, these algorithms are typically designed to achieve two goals: (1) balancing the exploration-exploitation trade-off, and (2) minimizing cumulative regret. In this section, we describe core algorithmic strategies for nonlinear contextual bandit problems, where the expected reward is modeled as a nonlinear function of the arms and, possibly, covariates. Below, we provide a general algorithmic template, followed by detailed descriptions of three model-based strategies, namely, $\epsilon$-greedy, Upper Confidence Bound (UCB), and ViOlin, highlighting how each balances exploration and exploitation. Note that all the three algorithms have been studied in various bandit problems before (Auer and Ortner,, 2010; Arya and Yang,, 2020; Dong et al.,, 2021), we propose model-based versions here.
3.1 General framework
Recall, $f(x;\theta)$ denotes a known nonlinear model (e.g., Michaelis-Menten, logistic, quadratic) parameterized by $\theta$, where $x\in\mathcal{X}$ is the action (e.g., N-rate) and the expected reward is $\mathbb{E}[Y\mid x]=f(x;\theta)$. At each time $t$, the agent selects an action $x_{t}$, then receives a reward $y_{t}$.
Algorithm 1 General Model-based Nonlinear Contextual Bandit Framework
1:Input: Action set $\mathcal{X}$, nonlinear model class $f(\cdot;\theta)$, time horizon: $T$
2:Initialize: Pull each arm at least once and receive corresponding rewards until round $t_{0}$
3:for $t=t_{0}+1$ to $T$ do
4:  Estimate model parameters $\hat{\theta}_{t}$ using past data $\{(x_{s},y_{s}):s<t\}$
5:  Choose action $x_{t}$ using a strategy/policy $\pi$ utilizing $f(x;\hat{\theta}_{t})$
6:  Observe reward $y_{t}$
7:  Update dataset with $(x_{t},y_{t})$
8:end for
In the algorithm sketch (Algorithm 1), in step 3-4, at each round $t$, the decision-maker uses all previous data to update its understanding of the nonlinear reward model by estimating the unknown parameters $\theta$ that describe the model. Given past measurements of arm-reward pairs, we estimate the model parameters using nonlinear least squares, e.g., via the curve_fit function in Python’s scipy.optimize package. For a model $f(x;\theta)$ with parameters $\theta$, we solve
$\hat{\theta}=\arg\min_{\theta}\sum_{i=1}^{n}\ell(y_{i},f(x_{i};\theta)),$
where $\ell$ is a loss function (such as squared-error loss) computed using the observed $(x_{i},y_{i})$ pairs, with the goal of minimizing the loss. This practical approach is central to our implementation, allowing bandit algorithms to leverage data-driven parameter estimates and domain-specific models. Then, in step 5, the decision-maker chooses the next action by balancing the exploration-exploitation trade-off and satisfying the criterion for cumulative reward maximization. Then, in step 6, the reward is observed after the season and is added to the dataset.
Remark 1.
Unlike standard supervised learning or a fully-factorial experiments, which provides full feedback by revealing the outcome for every possible action/arm, bandit algorithms operate under partial feedback: at each round, only the reward corresponding to the action actually chosen is observed. This is often referred to as bandit feedback.
Next, we describe three model-based algorithms that follow the general structure of Algorithm 1 but differ in how they balance the exploration-exploitation trade-off.
3.2 Epsilon-Greedy
The $\epsilon$-greedy algorithm (Algorithm 2) (Langford and Zhang,, 2008; Yang and Zhu,, 2002) is one of the simplest approaches for balancing exploration and exploitation. As the name suggests, most of the time it is greedy, i.e., it chooses the action that appears best according to current knowledge (exploitation), but with small probability $\epsilon$, it selects an action at random (exploration). This mechanism ensures continued learning, preventing the algorithm from becoming stuck on suboptimal choices.
Several variants of $\epsilon$-greedy exist; the most popular is the annealed or decaying $\epsilon$-greedy, where the exploration probability $\epsilon_{t}$ is a non-increasing sequence that tends to zero as $t\rightarrow\infty$. The intuition is that, as more data are collected and the model becomes more certain, the algorithm increasingly favors exploitation over exploration. Note that $\epsilon$ (or the sequence $(\epsilon_{t})$) is specified by the user, allowing the decision-maker to directly control the balance between exploration and exploitation; this choice directly influences the cumulative regret.
Epsilon-greedy is simple, robust, and easy to implement. It is particularly useful when the model is misspecified or when simple, interpretable algorithms are desired.
Algorithm 2 Epsilon-Greedy for Model-based Nonlinear Contextual Bandits
1:Input: Action set $\mathcal{X}$, Exploration parameter $\epsilon>0$, non-linear model class $f(\cdot;\theta)$
2:Initialize: Random pull arms and receive corresponding rewards until round $t_{0}$
3:for $t=t_{0}+1$ to $T$ do
4:  Estimate $\hat{\theta}_{t}$ from previous data: $\{x_{1},y_{1},\ldots,x_{t-1},y_{t-1}\}$
5:  Select action $x_{t}$ according to:
$x_{t}=\begin{cases}\text{sample uniformly from }\mathcal{X}&\text{with probability }\epsilon\\ \arg\max_{x\in\mathcal{X}}f(x;\hat{\theta}_{t})&\text{with probability }1-\epsilon\end{cases}$
6:  Observe $y_{t}$, update data
7:end for
If $\mathcal{X}$ is a finite set of $K$ elements, “sample uniformly from $\mathcal{X}$” means selecting an element at random with equal probability, so that each arm is chosen with probability $1/K$. In the $\epsilon$-greedy algorithm, during the exploration step, each arm is selected with probability $\epsilon/K$. If $\mathcal{X}$ is a continuous interval or region, this means sampling from the continuous uniform distribution over $\mathcal{X}$, i.e., $x_{t}\sim\mathrm{Unif}(\mathcal{X})$.
3.3 Nonlinear model-based UCB
Overview:
The UCB (Upper Confidence Bound) algorithm (Auer and Ortner,, 2010; Chu et al.,, 2011; Zhou et al.,, 2020) embodies the principle of “optimism in the face of uncertainty” by selecting actions that maximize an upper confidence bound on the expected reward. In each round, it favors actions that either appear promising based on current estimates or for which the uncertainty is high, thereby systematically encouraging exploration as well as exploitation.
In model-based nonlinear regression, the uncertainty $\mathrm{Unc}_{t}(x)$ is often based on the variability of the estimated reward due to finite data, as quantified via the standard error of the model prediction. Specifically, the uncertainty term is given by
$\mathrm{Unc}_{t}(x)=\sqrt{\nabla_{\theta}f(x;\hat{\theta}_{t})^{\top}\,\widehat{\mathrm{Cov}}(\hat{\theta}_{t})\,\nabla_{\theta}f(x;\hat{\theta}_{t})},$ (4)
where $\nabla_{\theta}f(x;\hat{\theta}_{t})$ denotes the gradient of the reward model with respect to the parameters, evaluated at the current estimate $\hat{\theta}_{t}$, and $\widehat{\mathrm{Cov}}(\hat{\theta}_{t})$ is the estimated covariance matrix of $\hat{\theta}_{t}$.
The UCB action selection rule then takes the form
$x_{t}=\arg\max_{x\in\mathcal{X}}\,f(x;\hat{\theta}_{t})+\alpha\cdot\mathrm{Unc}_{t}(x)$ (5)
where $\alpha>0$ is a user-specified confidence parameter that controls the degree of exploration.
UCB algorithms provide strong theoretical guarantees for balancing exploration and exploitation, and are particularly effective when reliable, model-based estimates of prediction uncertainty are available (see, e.g., (Zhou et al.,, 2020)). It is particularly advantageous in settings where data collection is expensive or risky, as it prioritizes actions that could yield high rewards or that have not been thoroughly explored.
Algorithm 3 UCB for Nonlinear Contextual Bandits
1:Input: Action set $\mathcal{X}$, confidence parameter $\alpha>0$
2:for $t=1$ to $T$ do
3:  Estimate $\hat{\theta}_{t}$ from the previous data
4:  for each $x\in\mathcal{X}$ do
5:   Compute predicted reward $f(x,z_{t};\hat{\theta}_{t})$ and uncertainty $\mathrm{Unc}_{t}(x)$ as defined in (4)
6:   Calculate $\text{UCB}_{t}(x)=f(x,z_{t};\hat{\theta}_{t})+\alpha\cdot\mathrm{Unc}_{t}(x)$
7:  end for
8:  Choose $x_{t}=\arg\max_{x\in\mathcal{X}}\text{UCB}_{t}(x)$
9:  Observe $y_{t}$, update data
10:end for
3.4 ViOlin (Virtual Ascent with Online Model Learner)
The ViOlin algorithm (Algorithm 4(Dong et al.,, 2021) is a model-based bandit strategy designed for efficient learning in nonlinear settings. Unlike $\epsilon$-greedy, which explores actions uniformly at random, or UCB, which prioritizes actions with high uncertainty, ViOlin is a greedy method that leverages both the current model estimate and local geometric information about the reward surface (i.e., gradient and curvature). At each round, the algorithm chooses the action predicted to be the best for profit maximization by the current estimated model, then it updates the model using not only the observed rewards, but also estimates of the local gradient and Hessian of the reward function at the selected action. This curvature-matching approach yields strong sample efficiency guarantees, especially in complex, high-dimensional, or highly nonlinear problems, by ensuring the model learns both the value and the local shape of the reward function. While originally, Dong et al., (2021) present the ViOlin algorithm for the non-noisy setting, we illustrate its performance under model noise as well.
Algorithm 4 ViOlin: Virtual Ascent with Online Model Learner (Dong et al.,, 2021)
1:Input: Model class $\mathcal{F}$ parameterized by $\theta$, action set $\mathcal{X}$, initial model $\hat{\theta}_{1}$, online learner $\mathcal{O}$, total rounds $T$
2:for $t=1$ to $T$ do
3:  Model Update: Update model parameters $\hat{\theta}_{t}$ using all data observed so far via the online learner $\mathcal{O}$
4:  Action Selection: Choose
$x_{t}=\arg\max_{x\in\mathcal{X}}f(x;\hat{\theta}_{t})$
5:  Play and Observe: Play action $x_{t}$, observe the true reward $r_{t}=f(\theta^{*},x_{t})+\epsilon_{t}$
6:  Curvature Estimation: Estimate the gradient and Hessian of the real reward function at $x_{t}$ using random perturbations or finite differences:
$\widehat{\nabla}_{x}f(\theta^{*},x_{t}),\quad\widehat{\nabla}^{2}_{x}f(\theta^{*},x_{t})$
7:  Curvature-Matching Model Update: Update $\mathcal{O}$ by minimizing
$\ell_{t}(\theta)=\;$ $(f(x_{t};\theta)-y_{t})^{2}+\alpha_{1}\left\|\nabla_{x}f(x_{t};\theta)-\widehat{\nabla}_{x}f^{*}(x_{t})\right\|^{2}$
$+\alpha_{2}\left\|\nabla^{2}_{x}f(x_{t};\theta)-\widehat{\nabla}^{2}_{x}f^{*}(x_{t})\right\|^{2}$
8:end for
ViOlin is particularly attractive in data-limited settings (e.g., smallholder or resource-constrained agricultural experiments), where minimizing the number of field trials is crucial. ViOlin’s greedy exploration is guided by the model fit and its local geometry, without explicit randomized or uncertainty-driven action selection. This enables fast convergence to (local) optima when the model class matches the true reward structure and curvature information is reliable. For well-behaved, low-dimensional agronomic models, the performance of ViOlin may be comparable to that of simpler methods such as $\epsilon$-greedy and UCB. However, as model complexity such as multi-modal response curves or as dimensionality of $\mathcal{X}$ increases, ViOlin’s sample efficiency guarantees make it especially advantageous. This makes ViOlin well suited for adaptive experimentation in limited data regimes, such as on-farm field trials or precision agriculture for smallholder contexts, where each data point is costly to obtain.
Summary of algorithm differences:
$\epsilon$-greedy explores uniformly at random, UCB targets actions with high uncertainty (optimism-based exploration), and ViOlin is greedy but uses strong model-based curvature matching for sample-efficient learning in complex settings.
While the primary emphasis of this work is not on theoretical development, we present and discuss several well-known theoretical results and highlight their implications for the nonlinear model-based contextual bandit framework.
4 Theoretical guarantees for nonlinear bandit algorithms
In this section, we characterize the regret performance of sequential algorithms under increasingly flexible model-based reward mechanisms, beginning with classical linear models, then considering nonparametric function classes, nonlinear parametric (mechanistic) models as used in agronomy, and finally accounting for the impact of model misspecification.
Minimax regret for linear and GLM bandits.
For linear bandits with $d$-dimensional parameterization, the minimax expected cumulative regret is known to satisfy
$\text{E}[R_{T}]=O\left(d\sqrt{T}\log T\right),$
where $T$ is the time horizon (Abbasi-Yadkori et al.,, 2011; Dani et al.,, 2008). Generalized linear bandit models achieve similar rates up to logarithmic factors (Filippi et al.,, 2010; Bastani et al.,, 2020).
Nonparametric bandits.
For more general nonparametric function classes, such as Lipschitz or RKHS (kernelized) reward functions, the minimax regret scales polynomially with the dimension $d$ of the action/context space (Slivkins,, 2014; Srinivas et al.,, 2012):
$R_{T}=\widetilde{O}\left(T^{\frac{d+1}{d+2}}\right),$
where $d$ is the relevant dimension. These models are highly flexible but suffer from poor sample efficiency in high dimensions, motivating the use of structured parametric models when such domain knowledge is available.
NeuralUCB and nonlinear function classes.
Recent advances have extended bandit algorithms to rich nonlinear function classes via neural networks. In particular, the NeuralUCB algorithm (Zhou et al.,, 2020) achieves a regret bound of
$R_{T}=\widetilde{O}\left(\tilde{d}(\lambda)\sqrt{T}\right),$
where $\tilde{d}(\lambda)$ denotes the effective dimension of the neural tangent kernel (NTK) associated with the underlying neural network, and $\widetilde{O}$ suppresses logarithmic dependencies on $T$ and $1/\lambda$. The effective dimension is defined as
$\tilde{d}(\lambda)=\mathrm{Tr}\left(K_{\mathcal{X},\mathcal{X}}(K_{\mathcal{X},\mathcal{X}}+\lambda I)^{-1}\right),$
where $K_{\mathcal{X},\mathcal{X}}$ is the NTK Gram matrix over the sequence of played arms $\{x_{1},\dots,x_{T}\}$ and $\lambda>0$ is a regularization parameter. For a two-layer ReLU neural network with input dimension $d$ and width $m$, it holds that $\tilde{d}(\lambda)\leq d\,\mathrm{poly}(m,\log(T/\lambda))$, so that the regret rate can be written as
$R_{T}=\widetilde{O}\left(d\,\mathrm{poly}(m)\sqrt{T}\right).$
Here, $\mathrm{poly}(m,\cdot)$ denotes a polynomial function of the network width and the indicated arguments. Thus, the regret bound increases with both the expressiveness of the neural network (through $m$) and the ambient input dimension $d$, but retains the $\sqrt{T}$ scaling characteristic of parametric bandit models. This result demonstrates that, under suitable overparameterization, nonlinear bandits with neural function approximation can achieve regret scaling close to the linear case (up to polynomial factors in $m$), substantially improving sample efficiency over generic nonparametric rates. However, this improvement is obtained at the cost of increased computational complexity and model tuning.
Table 1: Summary of Regret Rates for Contextual Bandits under Different Model Classes
Model Class Regret Bound Reference
Linear Bandit $O(d\sqrt{T}\log T)$ Abbasi-Yadkori et al., (2011)
GLM Bandit $\widetilde{O}(d\sqrt{T})$ Filippi et al., (2010)
Nonparametric (Lipschitz, RKHS) $\widetilde{O}(T^{\frac{d+1}{d+2}})$ Slivkins, (2014)
NeuralUCB (2-layer NN) $\widetilde{O}(d\,\mathrm{poly}(m)\sqrt{T})$ Zhou et al., (2020)
4.1 Model-based nonlinear bandits.
Building on these insights, we now focus on model-based nonlinear bandits, a regime particularly relevant in agronomy, where biological processes often admit interpretable, low-dimensional parameterizations. Since one of our motivations for this work is to provide decision-making algorithms in sample-limited scenarios, we focus on characterizing sample complexity more than providing regret guarantees in this section. Recent work by Dong et al., (2021) has shown a breakthrough for model-based nonlinear bandits: for any class of reward functions with bounded sequential Rademacher complexity, it is possible to find an $\epsilon$-approximate local maximum with sample complexity polynomial in the complexity of the model class, independent of the action dimension. This suggests that, unlike classical methods whose sample complexity or regret is exponential in dimension, model-based nonlinear bandits can be vastly more efficient when the reward function admits a suitable low-complexity parameterization.
Sample Complexity and Local Regret for Nonlinear Bandits.
In the nonlinear bandit setting, the goal is not always to find a global maximum (which is often computationally and statistically intractable), but rather to identify an approximate local maximum of the unknown reward function as efficiently as possible. Here, sample complexity refers to the minimum number of experimental rounds required to guarantee, with high probability, that the recommended action achieves performance within a specified tolerance of a local optimum. This guarantee is formalized through the notion of local regret, which measures the suboptimality of the chosen action relative to the best locally optimal action. In practice, for well-behaved (e.g., unimodal) functions such as those common in agronomic yield response, a local maximum often coincides with the global maximum.
In particular, a point $x$ is said to be an $(\epsilon_{g},\epsilon_{h})$-approximate local maximum if its gradient is at most $\epsilon_{g}$ (i.e., $\|\nabla f(x)\|_{2}\leq\epsilon_{g}$) and the Hessian’s largest eigenvalue is at most $\epsilon_{h}$ (i.e., $f$ is sufficiently concave around $x$). The sample complexity quantifies how quickly an algorithm can find such a point.
In this work, we focus on reward functions (e.g., quadratic-plateau, Michaelis-Menten, Mitscherlich, and logistic) that satisfy the regularity conditions (such as bounded gradients and Hessians) assumed in recent theoretical results (Dong et al.,, 2021). These properties ensure that our algorithms are well-behaved: the models are smooth enough for efficient learning, and the theoretical sample complexity guarantees apply. As a result, we can meaningfully compare algorithms in terms of how many rounds are needed to reach near-optimal fertilizer recommendations with high confidence. We first define the notion of sequential Rademacher complexity and then state Theorem 1.1 from Dong et al., (2021) in terms of sequential complexity, and then we tailor the result to our specific non-linear model classes.
Intuitively, sequential Rademacher complexity measures how hard it is for a learning algorithm to reliably make good decisions when facing an environment that can adapt to the algorithm’s past actions. More formally, the sequential Rademacher complexity (Rakhlin et al., 2015b, ) is defined as follows.
Definition 1 (Sequential Rademacher Complexity).
Let $\mathcal{F}$ be a class of real-valued functions defined on $\mathcal{X}$. The sequential Rademacher complexity of $\mathcal{F}$ over $T$ rounds is defined as
$\mathfrak{R}_{T}^{\mathrm{seq}}(\mathcal{F})=\sup_{x_{1},\ldots,x_{T}}\mathbb{E}_{\epsilon}\left[\sup_{f\in\mathcal{F}}\frac{1}{T}\sum_{t=1}^{T}\epsilon_{t}f\left(x_{t}(\epsilon_{1},\ldots,\epsilon_{t-1})\right)\right]$
where the supremum is taken over all sequences of functions $x_{t}:\{-1,+1\}^{t-1}\to\mathcal{X}$, and $\epsilon_{1},\dots,\epsilon_{T}$ are independent Rademacher random variables, i.e., $\mathbb{P}(\epsilon_{t}=1)=\mathbb{P}(\epsilon_{t}=-1)=\frac{1}{2}$ for each $t$.
Now, we present the result by Dong et al., (2021) that determines the sample complexity rate for ViOlin.
Theorem 1 (Sample Complexity for Model-based Nonlinear Bandits (Dong et al.,, 2021, Theorem 1.1)).
Suppose the sequential Rademacher complexity of the loss function class $\mathfrak{R}_{T}^{\mathrm{seq}}(\mathcal{F})$ induced by the reward function class $\{f(\theta,\cdot):\theta\in\Theta\}$ is bounded by $\sqrt{R(\Theta)T\,\mathrm{polylog}(T)}$ for some complexity parameter $R(\Theta)$. Then, there exists an algorithm (ViOlin)) (Algorithm  4) that finds an $\epsilon$-approximate local maximum with $\widetilde{O}(R(\Theta)\epsilon^{-8})$ samples, independent of the dimension of the action space.
Note that most of the agronomic models assume bounded non-linear functions, such as ones that exhibit plateauing behavior after a certain threshold. Therefore, it is of interest to quantify the sequential Rademacher complexity for these class of functions. For parametric classes with bounded functions and inputs (such as those considered in this paper), we prove a theoretical bound on the sequential Rademacher complexity of bounded functions and show that indeed for this class of functions the sequential Rademacher complexity has an upper bound as required in Theorem 1.
Theorem 2 (Sequential Rademacher Complexity for Bounded Functions).
Let $\mathcal{F}\subseteq[-B,B]^{\mathcal{D}}$ be a class of functions uniformly bounded by $B>0$. Then the sequential Rademacher complexity of $\mathcal{F}$ satisfies
$\mathfrak{R}_{T}^{\mathrm{seq}}(\mathcal{F}):=\sup_{\boldsymbol{x}}\mathbb{E}_{\boldsymbol{\epsilon}}\left[\sup_{f\in\mathcal{F}}\frac{1}{T}\sum_{t=1}^{T}\epsilon_{t}f(x_{t}(\epsilon_{1},\ldots,\epsilon_{t-1}))\right]\leq C\cdot B\cdot\sqrt{\frac{\log T}{T}},$
for some universal constant $C>0$.
We defer the proof of Theorem 2 to the Appendix (Section A) for brevity and discuss its implications from an application viewpoint. Our theoretical result complement Dong et al., (2021)’s general framework by concretely instantiating the sample complexity for these practically important nonlinear classes, including quadratic plateau, Michaelis-Menten, Mitscherlich, and logistic models—the sequential Rademacher complexity $\mathfrak{R}_{T}^{\mathrm{seq}}(\mathcal{F})$ can be bounded as $\widetilde{O}(1/\sqrt{T})$. Therefore, the sample complexity for identifying an $\epsilon$-optimal arm (in the sense of a local maximum) in these bandit settings scales as $\widetilde{O}(1/\epsilon^{8})$, with constants depending on the parameter and input bounds. Note that, this result of Dong et al., (2021) is not tight and can potentially be improved in terms of the polynomial dependence on $\epsilon$.
Corollary 1 (Sample Complexity for online learning of Bounded Non-linear Reward classes).
Using the bound in Theorem 2 in Theorem 1, an $\epsilon$-optimal solution (arm) can be identified with sample complexity (number of time steps), $\widetilde{O}(1/\epsilon^{8})$ for the ViOlin algorithm of (Dong et al.,, 2021), independent of the action (arm) space dimension.
Interpretation.
This shows that for commonly used nonlinear agronomic yield response models, bandit algorithms can efficiently identify near-optimal fertilizer rates with a number of field trials that scales polynomially with the desired accuracy and avoids the curse of dimensionality in the action space. Our results instantiate Dong et al., (2021)’s general theory for these specific models, providing practical sample complexity bounds for adaptive experimentation in agriculture. In Theorem 1, the parameter $\epsilon>0$ represents the maximum allowable gap between the algorithm’s selected action (e.g., a nitrogen rate recommendation) and the optimal (locally maximizing) action in terms of expected reward. That is, with high probability, the bandit algorithm identifies an action $x$ such that
$f^{*}(x^{*})-f^{*}(x)\leq\epsilon,$
where $f^{*}$ is the true (unknown) reward or profit function, $x^{*}$ is a (local) maximizer of $f^{*}$, and $x$ is the action recommended by the algorithm.
The sample complexity required to achieve an $\epsilon$-optimal recommendation scales polynomially with $1/\epsilon$ (specifically, as $O(1/\epsilon^{8})$ in Theorem 1), so tighter tolerances (smaller $\epsilon$) require proportionally more field trials. In practice, $\epsilon$ should be set according to what constitutes an economically meaningful margin for decision-making.
While Dong et al., (2021) establish general sample complexity and regret guarantees for nonlinear bandit problems for their proposed algorithm (Algorithm 4) including highly complex model classes like 2-layer Neural networks, our agricultural decision models are low-dimensional and interpretable (e.g., quadratic threshold). In this favorable case, classic algorithms such as UCB and $\epsilon$-greedy are not only interpretable and easy to implement but also benefit from even stronger theoretical guarantees than those stated in Dong et al., (2021), due to the reduced complexity of the model class. Therefore, we use these classic model-based approaches and provide sample complexity statements derived from Dong et al., (2021) as a conservative bound.
Under Model Misspecification
In many applications, including agricultural yield optimization, the true reward function may not be perfectly captured by any member of a chosen parametric model class (such as quadratic plateau or Michaelis-Menten). This is known as model misspecification, and can have important implications for the performance and guarantees of bandit algorithms.
Most contextual bandit algorithms rely on the assumption of a well-specified model (or “realizability”), that is, the existence of a function $f^{*}\in\mathcal{F}$ such that the mean reward is $\mu(x)=f^{*}(x)$ for all actions $x$ (Chu et al.,, 2011; Abbasi-Yadkori et al.,, 2011; Agarwal et al.,, 2012; Foster et al.,, 2018). However, in practice, exact realizability rarely holds, and it is important to consider the effect of model misspecification.
Recent work (see, e.g., (Greenewald et al.,, 2021; Lattimore et al.,, 2020; Ghosh et al.,, 2017)) has studied contextual bandits under various notions of misspecification. A common formulation is the uniform $\varepsilon$-misspecification setting, where
$\inf_{f\in\mathcal{F}}\sup_{x\in\mathcal{X}}\left|\mu(x)-f(x)\right|\leq\varepsilon,$ (6)
for some $\varepsilon>0$.
For linear contextual bandits, Lattimore et al., (2020) show that the cumulative regret must satisfy
$R_{T}\gtrsim d\sqrt{T}+\varepsilon\sqrt{d}T,$
where $d$ is the model dimension and $T$ is the time horizon. The first term corresponds to the minimax regret under well-specified models, while the second is an unavoidable “price of misspecification”. In our setting, the models (quadratic plateau, logistic, etc.) may be misspecified relative to the true reward. As such, the regret of our bandit algorithms can be interpreted as
$R_{T}=O(\text{model-based regret})+T\cdot\varepsilon_{T}(S),$
where the first term is the regret under realizability, and the second term is proportional to the sequence-specific misspecification error. This highlights the need for robust algorithms and motivates future work on model selection or hybrid model-based/nonparametric approaches.
5 Simulation study: Emulating real-world agricultural decision-making
5.1 From yield to profit: Economic optimization of fertilizer
In agricultural decision-making, maximizing yield alone does not necessarily lead to optimal outcomes. Most agronomic yield functions exhibit diminishing returns to fertilizer input, i.e., yield increases with fertilizer up to a point and then plateaus. Since fertilizer is a costly input, applying more of it beyond the agronomically efficient range leads to higher expenses without commensurate gains in yield. As a result, the economically optimal fertilizer dose is not the one that produces the maximum yield, but rather the one that maximizes profit, the difference between revenue and input cost.
To formalize this, let $p_{y}$ denote the price per unit yield (e.g., \$/bushel), and $p_{x}$ the price per unit of fertilizer (e.g., \$/lb N). Suppose the yield response is described by a parametric function $f(x;\theta)$, where $x$ is the fertilizer application level and $\theta$ denotes model parameters capturing biological or agronomic characteristics. The corresponding economic profit is then given by
$\Pi(x)=p_{y}\cdot f(x;\theta)-p_{x}\cdot x.$ (7)
The economically optimal fertilizer level $x^{*}$ maximizes this profit:
$x^{*}=\arg\max_{x\in\mathcal{X}}\,\Pi(x)=\arg\max_{x\in\mathcal{X}}\,\left[p_{y}f(x;\theta)-p_{x}x\right],$ (8)
where $\mathcal{X}$ is the feasible set of fertilizer rates (e.g., a grid from 0 to 250 lb N/ac). Depending on the form of $f$, this maximization may admit a closed-form solution or may require numerical optimization. These generic routines can accommodate any parametric or nonparametric yield response function $f(x;\theta)$, enabling interpretable economic optimization across a broad range of agronomic scenarios. For the nonlinear yield functions studied in our simulations, closed-form expressions for $x^{*}$ can be derived and are summarized in Table 2. These expressions highlight how the optimal fertilizer rate depends jointly on agronomic parameters $\theta$ and economic parameters $p_{y},p_{x}$, all of which must be estimated from data.
We now turn to the problem of sequentially optimizing fertilizer decisions using model-based bandit algorithms that account for uncertainty and limited data. At each round, the algorithm fits a nonlinear yield response model to the observed data and selects the fertilizer rate that maximizes the estimated economic profit, optionally incorporating an uncertainty-based exploration bonus. We investigate three nonlinear model-based strategies:
  1. 1.
    Model-based $\epsilon$-greedy algorithm (Algorithm 5): This method selects a random fertilizer rate with probability $\epsilon_{t}=t^{-a}$, with $a$ chosen from $\{0.5,1,1.5\}$, for exploration. Otherwise, it chooses the rate closest to the current profit-maximizing estimate $x^{*}$, computed using the fitted model and closed-form expressions where available.
  2. 2.
    Model-based UCB algorithm (Algorithm 6): This approach augments the estimated profit at each arm with a model-based uncertainty term, derived from the parameter covariance of the fitted nonlinear model, and selects the rate with the highest upper confidence bound.
  3. 3.
    ViOlin) algorithm (Algorithm 4): A curvature-aware strategy that exploits local second-order structure in the profit function to accelerate convergence, particularly under smooth nonlinearity.
Although these methods differ in their exploration strategies, they all share the same structure: fitting a nonlinear model to guide adaptive decision-making under uncertainty. Accurate initialization of nonlinear model parameters is crucial to the success of these algorithms, especially in small-sample settings. Poor initial values can mislead early optimization steps and delay convergence. To benchmark the performance of nonlinear model-based strategies, we compare them against two widely used baselines: LinUCB (Chu et al.,, 2011) and $k$-Nearest Neighbor-UCB (kNN-UCB) (Reeve et al.,, 2018).
  1. 4.
    LinUCB algorithm (Algorithm 7): A linear bandit method that models yield as a linear function of fertilizer rate and selects actions using upper confidence bounds derived from online regression.
  2. 5.
    kNN-UCB algorithm (Algorithm 8): A nonparametric method that estimates yield using the $k$ nearest neighbors of each candidate arm and quantifies uncertainty via local sample variance.
Further descriptions of these algorithms can be found in the Appendix. These baselines span the spectrum from parametric simplicity to nonparametric flexibility and help highlight the advantages of domain-informed nonlinear models in sample-constrained settings.
Table 2: Closed-form expressions for the profit-maximizing fertilizer dose $x^{*}$ for different yield response models.
Model Yield Function $Y(x)$ Profit-maximizing $x^{*}$
Mitscherlich $A(1-e^{-bx})$ $\frac{1}{b}\ln\left(\frac{p_{x}}{p_{y}Ab}\right)$
Quadratic (threshold) $\begin{array}[]{l}a+bx+cx^{2},\ x\leq x_{0}\\ a+bx_{0}+cx_{0}^{2},\ x>x_{0}\end{array}$ $\min\left\{x_{0},\ \max\left\{0,\ \frac{1}{2c}\left(\frac{p_{x}}{p_{y}}-b\right)\right\}\right\}$
Michaelis-Menten $\dfrac{ax}{b+x}$ $\sqrt{\dfrac{abp_{y}}{p_{x}}}-b$
Logistic $\dfrac{A}{1+e^{-B(x-C)}}$ $\begin{array}[]{l}u^{*}=\dfrac{\gamma-2-\sqrt{(\gamma-2)^{2}-4}}{2},\\ \gamma=\dfrac{Bp_{y}A}{p_{x}},\quad x^{*}=C-\dfrac{1}{B}\ln u^{*}\end{array}$
Algorithm 5 Model-Based $\epsilon$-Greedy for Economic Profit Maximization
0: Fertilizer levels $\mathcal{X}$, time horizon $T$, prices $p_{y},p_{x}$, exploration rate $\epsilon_{t}$
1: Initialize dataset $\mathcal{D}\leftarrow\emptyset$
2:for $t=1$ to $T$ do
3:  if $u\sim\text{Uniform}(0,1)<\epsilon_{t}$ or $|\mathcal{D}|<n_{0}$ then
4:   Select fertilizer level $x_{t}\sim\text{Uniform}(\mathcal{X})$ // Explore
5:  else
[…]
7:   Compute $x^{*}=\arg\max_{x\in\mathcal{X}}\,\left[p_{y}f(x;\hat{\theta})-p_{x}x\right]$
8:   Select $x_{t}\in\mathcal{X}$ closest to $x^{*}$
9:  end if
10:  Apply $x_{t}$, observe yield $Y_{t}$
11:  Compute profit: $\Pi_{t}=p_{y}\cdot Y_{t}-p_{x}\cdot x_{t}$
12:  Add $(x_{t},Y_{t})$ to dataset $\mathcal{D}$
13:end for
Algorithm 6 Model-Based UCB for Economic Profit Maximization
0: Fertilizer levels $\mathcal{X}$, time horizon $T$, prices $p_{y},p_{x}$, UCB constant $\alpha$
1: Initialize dataset $\mathcal{D}\leftarrow\emptyset$
2:for $t=1$ to $T$ do
[…]
5:  for each $x\in\mathcal{X}$ do
6:   Predict profit: $\hat{\Pi}(x)=p_{y}\cdot f(x;\hat{\theta})-p_{x}\cdot x$
7:   Compute model-based uncertainty:
$\mathrm{Unc}(x)=p_{y}\cdot\sqrt{\nabla_{\theta}f(x;\hat{\theta})^{\top}\widehat{\mathrm{Cov}}(\hat{\theta})\nabla_{\theta}f(x;\hat{\theta})}$
8:   Compute UCB score: $UCB(x)=\hat{\Pi}(x)+\alpha\cdot\mathrm{Unc}(x)$
9:  end for
10:  Select $x_{t}=\arg\max_{x\in\mathcal{X}}UCB(x)$
11:  Apply $x_{t}$, observe yield $Y_{t}$
12:  Compute profit: $\Pi_{t}=p_{y}\cdot Y_{t}-p_{x}\cdot x_{t}$
13:  Update $\mathcal{D}\leftarrow\mathcal{D}\cup\{(x_{t},Y_{t})\}$
14:end for
Algorithm 7 LinUCB for Economic Profit Maximization
0: Fertilizer levels $\mathcal{X}$, time horizon $T$, prices $p_{y},p_{x}$, UCB parameter $\alpha$
1: Initialize dataset $\mathcal{D}\leftarrow\emptyset$
2: Define feature mapping $\varphi(x)=[1,\,x]^{\top}$
3:for $t=1$ to $T$ do
4:  if not enough data then
5:   Choose $x_{t}$ uniformly at random from $\mathcal{X}$
6:  else
7:   Fit linear model to data: $Y_{i}\approx\varphi(x_{i})^{\top}\hat{\beta}$
8:   For each $x\in\mathcal{X}$:
  • Predict yield: $\widehat{Y}(x)=\varphi(x)^{\top}\hat{\beta}$
  • Compute profit: $\hat{\Pi}(x)=p_{y}\cdot\widehat{Y}(x)-p_{x}\cdot x$
  • Estimate variance: $s^{2}(x)=\varphi(x)^{\top}(\mathbf{V}^{-1})\varphi(x)$, where $\mathbf{V}$ is the feature covariance matrix
  • Compute UCB score: $UCB(x)=\hat{\Pi}(x)+\alpha\cdot s(x)$
9:   Select $x_{t}=\arg\max_{x\in\mathcal{X}}UCB(x)$
10:  end if
11:  Apply $x_{t}$, observe $Y_{t}$
12:  Compute profit: $\Pi_{t}=p_{y}\cdot Y_{t}-p_{x}\cdot x_{t}$
13:  Append $(x_{t},Y_{t})$ to $\mathcal{D}$
14:end for
Algorithm 8 kNN-UCB for Economic Profit Maximization
0: Fertilizer levels $\mathcal{X}$, time horizon $T$, prices $p_{y},p_{x}$, UCB parameter $\alpha$, no. of neighbors $k$
1: Initialize dataset $\mathcal{D}\leftarrow\emptyset$
2:for $t=1$ to $T$ do
3:  if not enough data then
4:   Choose $x_{t}$ uniformly at random from $\mathcal{X}$
5:  else
6:   for each $x\in\mathcal{X}$ do
7:    Find $k$ nearest previously tried fertilizer rates to $x$ in $\mathcal{D}$
8:    Compute average yield: $\widehat{Y}_{k}(x)=\frac{1}{k}\sum_{j=1}^{k}Y_{j}$
9:    Compute sample standard deviation: $s_{k}(x)$ of yields among neighbors
10:    Compute profit: $\hat{\Pi}_{k}(x)=p_{y}\cdot\widehat{Y}_{k}(x)-p_{x}\cdot x$
11:    Compute UCB score: $UCB(x)=\hat{\Pi}_{k}(x)+\alpha\cdot\frac{s_{k}(x)}{\sqrt{k}}$
12:   end for
13:   Select $x_{t}=\arg\max_{x\in\mathcal{X}}UCB(x)$
14:  end if
15:  Apply $x_{t}$, observe $Y_{t}$
16:  Compute profit: $\Pi_{t}=p_{y}\cdot Y_{t}-p_{x}\cdot x_{t}$
17:  Append $(x_{t},Y_{t})$ to $\mathcal{D}$
18:end for
In the next section, we present simulation studies under both well-specified and misspecified conditions to compare the performance of these five bandit algorithms in terms of profit and cumulative regret.
5.2 Well-specified setting
Simulation setup.
In the well-specified setting, data were generated according to a quadratic threshold model of the form:
$Y(x)=\begin{cases}a+bx+cx^{2},&\text{if }x\leq x_{0}\\ a+bx_{0}+cx_{0}^{2},&\text{if }x>x_{0}\end{cases},$
where $Y(x)$ denotes crop yield (bu/ac) at fertilizer rate $x$ (lb N/ac), and $a=80$, $b=1.2$, c=0.003c=-0.003, $x_{0}=180$ were chosen to reflect realistic agronomic response curves. The profit at each round is computed as $\Pi(x)=p_{y}Y(x)-p_{x}x$, with grain price $p_{y}=\$5.00$ per bushel and fertilizer price $p_{x}\in\{0.3,0.5,0.7\}$ \$/lb N to mimic realistic prices per unit nitrogen fertilizer (such as Urea) in the Midwest US. Fertilizer rates are restricted to the discrete grid $\mathcal{X}=\{0,50,100,\ldots,250\}$ lb N/ac, and yields are corrupted by i.i.d. Gaussian noise with standard deviation $\sigma=0.5$.
We evaluated five algorithms: model-based $\epsilon$-greedy, nonlinear UCB, ViOlin (curvature-matched nonlinear bandit), LinUCB (linear model-based UCB), and kNN-UCB (nonparametric UCB using $k=3$ nearest neighbors). Hyperparameters for each of the algorithms are chosen based on theoretical guidelines and grid searches over multiple runs. We choose $\epsilon_{t}=t^{-0.5}$ for $\epsilon$-greedy in this setting, $\alpha=1$ in UCB, linUCB and kNN-UCB for uniformity and fairness in comparison, $\alpha_{1}=2.0,\alpha_{2}=640$ for the ViOlin algorithm based on guidelines in Dong et al., (2021). For each algorithm and parameter configuration, we performed $10$ independent simulation replicates of $T=30$ rounds. Each algorithm fits its specified model to the observed data and selects actions to maximize estimated profit. The regret at each round is defined as the difference between the average profit for the optimal arm (computed using knowledge of the true reward function) and the average profit realized by the arm selected by the algorithm, that is,
$\mathrm{Regret}_{t}=\mathbb{E}\left[\Pi(x^{*};\theta^{*})-\Pi(x_{t};\theta^{*})\right],$
where $x^{*}$ denotes the arm maximizing the true profit function, $\theta^{*}$ the true model parameters and $x_{t}$ is the arm selected at round $t$ by the bandit algorithm employed to make decisions. For each run, we recorded cumulative regret, average profit, and selected fertilizer rates.
Results and Interpretation.
Figure 1 presents the average cumulative regret and the average profit per round for each algorithm. Several trends are immediately apparent:
  • Superiority of nonlinear model-based methods: Both nonparametric (kNN-UCB) and linear (LinUCB) methods incur substantially higher regret than the nonlinear model-based approaches. In particular, kNN-UCB and LinUCB accumulate regret rapidly and fail to converge to near-optimal recommendations within the small sample budget. LinUCB, although it does not experience a sharp initial dip in profit due to exploration, tends to settle on a suboptimal arm and consistently incurs losses relative to the optimal strategy.
  • Exploration versus exploitation trade-off: All algorithms that actively explore—notably $\epsilon$-greedy and UCB—show a pronounced dip in profit at early rounds, reflecting the cost of exploratory actions. In contrast, the ViOlin algorithm, which is designed to be more exploitative (greedily maximizing predicted profit with curvature matching), avoids this initial dip and achieves high profits almost immediately. However, both $\epsilon$-greedy and UCB eventually recover and converge to policies with low cumulative regret, validating the effectiveness of their exploration in learning the optimal fertilizer rate.
Similar trends were observed across other values of the fertilizer price $p_{x}$ and for all nonlinear response models considered. Taken together, these simulation studies demonstrate a consistent advantage of model-based nonlinear bandit algorithms over both nonparametric and linear parametric alternatives in small-sample regimes. Based on these findings, we subsequently focus our comparisons on the family of nonlinear model-based algorithms, investigating their relative strengths under a range of simulation scenarios to clarify which approaches are most advantageous in which settings.
Figure 1: Comparing different Bandit algorithms for profit maximization/regret minimization
(a) Refer to caption
(b)Refer to caption
Figure 2: Selected fertilizer rates over time for the three non-linear model based bandit algorithms
In this next part of the simulation study, we focus on illustrating the implications of using the three non-linear model-based bandit algorithms for learning fertilizer rates sequentially with the goal of profit maximization. Again in a small-sample regime, we run each of our algorithms for $T=30$ rounds and replicate each run 10 times. Since our algorithms choose arms over time, it is important to visualize how the arm choices evolve over time for each of the algorithms. In Figure 2(a), we plot the running proportion of fertilizer rate ranging from $\{0,50,\ldots,250\}$ lb N/ac, selected at each round by each of the algorithm over the 10 replicates for the quadratic plateau model for $p_{x}=\$0.5$/lb N to mimic realistic prices per unit nitrogen fertilizer (Urea) in the Midwest US. Note that in the beginning, the UCB and $\epsilon$-greedy ($\epsilon_{t}=t^{-1.5}$) explore other arms such as $x=100,150$ lb N/ac, but just after a few rounds of exploration, learn the better arm to be $x=200$ lb N/ac. ViOlin is more greedy from the beginning itself as can be seen by the proportions of other arms (other than purple) selected in the initial rounds. In Figure 2(b), in order to assess the affect of increasing fertilizer prices on decision-making for profit maximization, we also plotted the average selected fertilizer rate over time for the three algorithms. We note that the optimum arm (fertilizer rate) choice decreases as $p_{x}$ (price per unit) increases from \$0.3 to \$0.7/lb N. Additionally, since our decisions depend significantly on how well our parameters for the non-linear model are estimated over time, in Figure 3, we also plot the parameter trajectories over time for the Quadratic Threshold model with $p_{X}=\$0.5$/lb N. Note that all the three algorithms, as data accumulates, the parameter estimates stabilize close to the true parameter values that were used to generate the data. In terms of interpretability, one can, for example, look at the parameter estimates for $x_{0}$ (bottom right), which denotes the threshold value at which the quadratic model plateaus, thus reflecting the saturating point over rounds (or seasons), beyond which additional fertilizer amount provide little benefit. We also checked for sensitivity to initial values and we noticed that good initial values matter for better convergence to the true parameters, especially in a small sample regime. Therefore, we recommend using domain-knowledge to set initial values of the parameters in the least-squares estimation algorithm. We conducted similar experiments for the other non-linear models and those results are presented in the Appendix.
Figure 3: Parameter trajectories for the quadratic threshold model over rounds
5.3 Misspecified setting
Simulation Setup.
To mimic a realistic situation in which the decision-maker fits an approximate but shape-compatible model, we generated data from the Mitscherlich response
$Y_{\mathrm{true}}(x)=A\!\left(1-e^{-b(x-d)}\right),\qquad A=120,\;b=0.015,\;d=80,$
yet fitted a quadratic-plateau curve
$Y_{\mathrm{fit}}(x)=a+bx+cx^{2}\;\text{for }x\leq x_{0},$
truncated at $x_{0}=180$ and takes the same value $a+bx_{0}+cx_{0}^{2}$ for $x>x_{0}$. All other ingredients are identical to the well-specified experiment except that we increase the time horizon to $T=100$, specifically, price grid $p_{x}\in\{0.3,0.5,0.7\}$ \$ lb-1 N, yield price $p_{y}=\$5.00$ bu-1, action set $\mathcal{X}=\{0,50,100,\ldots,250\}$ lb N ac-1, noise $\sigma=0.5$, and the 5 algorithms: $\epsilon$-greedy, nonlinear-UCB, ViOlin, LinUCB, kNN-UCB. Ten independent replicates were run for each algorithm and price.
Figure 4: Comparing different Bandit algorithms for cumulative regret over time in the misspecified setting.
(a) Refer to caption
(b)Refer to caption
Figure 5: Selected fertilizer rates over time for the three non-linear model based bandit algorithms for the misspecified setting
Figure 4 plots the average cumulative regret when the fitted quadratic-plateau and Michaelis Menten model is misspecified with respect to the true Mitscherlich process. Relative to the well-specified case, the model-based regret curves shift upwards, reflecting the price of fitting an approximate response surface. Yet their slopes remain moderate, indicating that the fitted quadratic-plateau still guides the search toward profitable regions quickly. Note, that similar to the well-specified setting, the model-based non-linear bandits perform better than the linear and non-parametric baselines. Although the quadratic-plateau form cannot capture the exponential approach to an asymptote exactly, its concave-then-flat shape matches the broad geometry of the Mitscherlich curve. As a consequence, all three nonlinear model-based algorithms, $\epsilon$-greedy, UCB, and ViOlin continue to accumulate substantially less regret than the linear (LinUCB) or non-parametric (kNN-UCB) baselines. This shape-compatible misspecification warrantees a more structured theoretical and experimental research, but our preliminary results are promising in this direction. Although ViOlin’s greedier, curvature-penalized score again avoids the large initial dip seen in UCB and $\epsilon$-greedy, note that UCB and $\epsilon$-greedy seem to perform comparably, especially in the Michaelis Menten fit, perhaps suggesting that exploration might help in misspecified settings. Other details can be found in the Appendix.
Figure 6: Parameter estimate trajectories for the misspecified setting
Similar to the well-specified setting, we also plot the running proportion of arms over time and the most frequently selected arm trajectories in Figure 4 (a) and (b) respectively. Although the arm selections remain mostly similar as in the well-specified settings, we do notice that UCB tends to explore more in the misspecified setting (with the same exploration parameter $\alpha=1$ as in the well-specified setting), which also results in better parameter estimates for UCB as is seen in Figure 6.
These findings underline a pragmatic principle for on-farm experimentation: when the shape of the assumed response is qualitatively correct, e.g. monotone-increasing with a plateau, model-based bandits remain sample efficient even if the functional form is wrong. Conversely, generic linear or purely non-parametric methods may demand far more data before converging, a luxury seldom available to smallholders. Hence, coupling modest domain knowledge with sequential decision-making promises a robust path toward Sustainable Development Goal 2 by enabling data-scarce producers to improve input efficiency without costly large-scale trials.
6 Conclusion and future work
In this work, we developed and systematically evaluated model-based nonlinear bandit algorithms for optimizing input decisions in agricultural systems. By framing the nitrogen rate optimization task as a sequential decision problem, we demonstrated that mechanistic model-based approaches, such as those relying on Mitscherlich, quadratic plateau, Michaelis-Menten, or logistic dose-response curves, can be flexibly integrated into the bandit framework to yield interpretable and effective decision-making strategies.
Our extensive simulation study revealed two principal findings. First, when the fitted model matches the true data-generating process, model-based bandit algorithms achieve substantially lower regret than nonparametric and linear baselines, highlighting the value of mechanistic knowledge. Second, when the fitted model is misspecified relative to the true reward structure, we observed a notable increase in regret, with the degree of performance loss depending on the extent of model mismatch. Notably, even under moderate misspecification, model-based bandits can still outperform linear methods such as LinUCB or fully model-free methods such as kNN-UCB, particularly when the mechanistic model provides a reasonable approximation of the underlying process and there are limited samples to learn from.
A key theoretical contribution of this work is the establishment of a conservative sample complexity bound for the ViOlin algorithm, which guarantees that the sample complexity, i.e., the number of rounds required to learn the local maxima of a non-linear reward function, scales polynomially in the desired accuracy and is independent of the dimension of the arm space. Although this result could be conservative for simpler models like we have considered, such applications further appeal to developing theoretical and methodological research that focuses on sample efficiency along with regret minimization. This is especially important in real-world applications like agriculture, where decision-makers often face tight resource constraints and require efficient learning algorithms that make the most out of limited data.
Our results underscore the importance of model selection and the trade-off between interpretability and robustness to misspecification in real-world sequential decision-making. Furthermore, we emphasize the indispensable role of domain experts in guiding data-driven algorithms, both by providing informed priors and initial conditions and by aiding in the interpretation of estimated parameters. Such collaborations ensure that adaptive management strategies remain both scientifically grounded and practically relevant.
Looking forward, several promising directions remain for future research. First, a systematic investigation into model selection or model averaging within the bandit setting could mitigate the risk of misspecification and enhance robustness. Extending the current framework to multi-dimensional or spatially varying decisions (i.e.,varying nitrogen rates across heterogeneous fields) represents an important direction, particularly given the prevalence of spatial heterogeneity in agricultural systems. Another practical extension is to incorporate weather and soil conditions as covariates in the non-linear models as these can play a significant role in the management decisions being made. Related to this is the fact that the non-linear models could be non-stationary and evolve over time. Developing strategies that adapt to non-stationarity in small sample settings is another relevant direction for future work. Another compelling direction is to bridge mechanistic and machine learning approaches, such as integrating semi-parametric or hybrid models that combine the interpretability of domain knowledge with the flexibility of data-driven components. Further, establishing tighter theoretical guarantees for regret under shape-constrained model misspecification, and extending these results to settings with delayed, censored, or partially observed rewards would further strengthen the practical applicability of bandit algorithms in complex, real-world domains.
From a broader perspective, the development of sample-efficient, interpretable, and robust sequential decision-making algorithms is crucial for advancing precision agriculture and achieving the Sustainable Development Goals (SDGs). These methods are particularly impactful for smallholder farmers operating with limited data, as they enable principled data-driven decision-making even under uncertainty and resource constraints. Overall, our study lays the groundwork for principled, adaptive management strategies in agriculture and beyond, with the potential for significant societal and environmental impact.
Disclosure Statement
The authors declare that there are no conflicts of interest regarding the publication of this paper.
Appendix
Appendix A Rademacher complexity for parametric nonlinear function classes
Proof of Theorem 2.
We follow the approach of Rakhlin et al., 2015b . Let $N_{1}(\alpha,\mathcal{F},T)$ denote the sequential covering number in the $\ell_{1}$-norm at scale $\alpha$. From equation (9) of their paper, we have:
$\mathfrak{R}_{T}(\mathcal{F})\leq\alpha+\sqrt{\frac{2\log N_{1}(\alpha,\mathcal{F},T)}{T}}.$ (9)
Using Corollary 1 from the same paper, the $\ell_{\infty}$-sequential covering number satisfies:
$N_{\infty}(\alpha,\mathcal{F},T)\leq\left(\frac{2eT}{\alpha}\right)^{\mathrm{fat}_{\alpha}(\mathcal{F})},$
and since $N_{1}(\alpha,\mathcal{F},T)\leq N_{\infty}(\alpha,\mathcal{F},T)$, we obtain:
$\log N_{1}(\alpha,\mathcal{F},T)\leq\mathrm{fat}_{\alpha}(\mathcal{F})\cdot\log\left(\frac{2eT}{\alpha}\right).$ (10)
For any class $\mathcal{F}\subseteq[-B,B]^{\mathcal{X}}$, we have from Theorem 10 of Rakhlin et al., 2015a :
$\mathrm{fat}_{\alpha}(\mathcal{F})\leq\left\lceil\frac{2B}{\alpha}\right\rceil.$ (11)
Now, using the bounds in (10) and (11) in (9) and choosing $\alpha=\frac{B}{\sqrt{T}}$, we get:
$\left\lceil\frac{2B}{\alpha}\right\rceil=\left\lceil 2\sqrt{T}\right\rceil\leq 3\sqrt{T},\quad\log\left(\frac{2eT}{\alpha}\right)=\log\left(\frac{2eT^{3/2}}{B}\right)=\mathcal{O}(\log T).$
Thus,
$\mathfrak{R}_{T}(\mathcal{F})\leq\frac{B}{\sqrt{T}}+\sqrt{\frac{6\sqrt{T}\cdot\log T}{T}}=\frac{B}{\sqrt{T}}+\sqrt{\frac{6\log T}{\sqrt{T}}}.$
The dominant term is $\frac{B\sqrt{\log T}}{\sqrt{T}}$, so we conclude:
$\mathfrak{R}_{T}(\mathcal{F})\leq C\cdot B\cdot\sqrt{\frac{\log T}{T}}$
for some universal constant $C$. ∎
Appendix B Finding optimal profit maximizing arms for different non-linear models
Recall that $\Pi(x)$ expresses the trade-off between revenue from yield and the cost of input. To find the economically optimal fertilizer dose $x^{*}$, we differentiate the profit function with respect to $x$: Below, we derive $x^{*}$ for each model considered in this work.
Mitscherlich Model:
The Mitscherlich yield response is
$Y(x)=A(1-e^{-bx}),$
so the profit function is
$\Pi(x)=p_{y}\,A(1-e^{-bx})-p_{x}x.$
The optimal dose $x^{*}$ solves
$\frac{d\Pi}{dx}=p_{y}\,A\,b\,e^{-bx}-p_{x}=0,$
giving
$x^{*}=-\frac{1}{b}\ln\left(\frac{p_{x}}{p_{y}Ab}\right).$
Note: If $\frac{p_{x}}{p_{y}Ab}\geq 1$, then $x^{*}=0$.
Quadratic Model with Threshold
Suppose
$Y(x)=\begin{cases}a+bx+cx^{2},&x\leq x_{0}\\ a+bx_{0}+cx_{0}^{2},&x>x_{0}\end{cases}$
with threshold $x_{0}$. Then
$\Pi(x)=p_{y}Y(x)-p_{x}x.$
For $x\leq x_{0}$, set the derivative to zero:
$\frac{d\Pi}{dx}=p_{y}(b+2cx)-p_{x}=0\implies x^{*}=\frac{1}{2c}\left(\frac{p_{x}}{p_{y}}-b\right).$
Thus,
$x^{*}=\min\left\{x_{0},\,\max\left\{0,\,\frac{1}{2c}\left(\frac{p_{x}}{p_{y}}-b\right)\right\}\right\}.$
If $x>x_{0}$, profit decreases with increasing $x$.
Michaelis-Menten (MM) Model
With
$Y(x)=\frac{ax}{b+x},$
the profit is
$\Pi(x)=p_{y}\,\frac{ax}{b+x}-p_{x}x.$
Setting the derivative to zero,
$\frac{d\Pi}{dx}=p_{y}\,\frac{ab}{(b+x)^{2}}-p_{x}=0\implies(b+x)^{2}=\frac{abp_{y}}{p_{x}}$
$x^{*}=\sqrt{\frac{abp_{y}}{p_{x}}}-b.$
We set $x^{*}=0$ if the right side is negative.
Logistic Model
For the logistic response
$Y(x)=\frac{A}{1+e^{-B(x-C)}},$
profit is
$\Pi(x)=p_{y}\,\frac{A}{1+e^{-B(x-C)}}-p_{x}x.$
Let $u=e^{-B(x-C)}$, so $x=C-\frac{1}{B}\ln u$. Setting the derivative to zero,
$p_{y}\,A\,\frac{Bu}{(1+u)^{2}}=p_{x},$
which yields a quadratic in $u$:
$u^{2}+\left(2-\frac{Bp_{y}A}{p_{x}}\right)u+1=0.$
Let
$\gamma=\frac{Bp_{y}A}{p_{x}}.$
The positive root is
$u^{*}=\frac{\gamma-2-\sqrt{(\gamma-2)^{2}-4}}{2},$
and the optimal dose is
$x^{*}=C-\frac{1}{B}\ln u^{*}.$
If $u^{*}\leq 0$ or $x^{*}<0$, set $x^{*}=0$.
Appendix C Baseline Algorithm descriptions
We compare the non-linear model based algorithms to the nonparametric and linear baselines. Specifically, we compare these with the LinUCB and $k$-Nearest Neighbor UCB algorithm. For the sake of completeness, we describe these algorithms below.
The LinUCB algorithm (Algorithm 7) is a classic model-based bandit method that assumes a linear relationship between the reward and the action (or its features). At each round, LinUCB fits a linear regression model to the observed data (line 2) and predicts the expected yield for each fertilizer rate. It computes an upper confidence bound (UCB) for each action by adding a model-based uncertainty term, proportional to the standard error of the prediction, to the profit estimate (lines 4–5). The fertilizer rate with the highest UCB is selected for the next trial (line 6), balancing exploration of uncertain arms and exploitation of high-yield arms. While LinUCB is efficient and easy to implement, its performance can degrade if the true reward function is nonlinear or misspecified.
The kNN-UCB algorithm (Algorithm 8) is a nonparametric bandit method that does not assume a specific model for the yield response. Instead, for each candidate fertilizer rate, it estimates the expected yield by averaging the observed yields of the $k$ most similar (nearest) rates previously tried (line 4). The uncertainty in this prediction is quantified by the sample variance among these $k$ neighbors (line 5). As with LinUCB, an upper confidence bound is constructed for each candidate rate (line 6), and the rate with the highest UCB is selected (line 7). Although kNN-UCB is robust to model misspecification and can capture complex nonlinearities, it generally requires more data to achieve accurate estimates, particularly when the action space is large or sparsely explored.
Appendix D Additional Simulation Results
In this section, we present the simulation results for other non-linear models that we considered in the simulation experiment of Section 5. First for the well-specified setting, we present the results for the Logistic and Mitscherlich model that was fit instead of Quadratic Plateau model. We consider the following true parameters for the simulation settings for each of the three models:
  • Logistic: $A=120,B=0.05,C=125,d=70$
  • Mitscherlich: $A=120,b=0.015,d=80$
Well-specified setting: Additional results
The same hyperparameter choices were made for the algorithms as in the well-specified setting in Section 5. Namely, we use $\epsilon_{t}=t^{-1.5}$ to encourage exploitation sooner than later, $\alpha=1$ for UCB and $\alpha_{1}=2,\alpha_{2}=640$ for ViOlin. Each algorithm was run 10 times and $T=30$ to emulate data-limited setting.
Figure 7: Cumulative Regret curves for the Logistic and Mitscherlich model.
(a) Refer to caption
(b)Refer to caption
Figure 8: Selected fertilizer rates over time for the three non-linear model based bandit algorithms for the well-specified setting corresponding to Logistic and Mitscherlich models, respectively. Note that the choice of arms changes based on the underlying true model.
It can be noted from Figure 8 that while the arm choices remain mostly similar in the Logistic model to that of Quadratic plateau model, for the Mitscherlich model the ViOlin algorithm seems to choose $x=250$ ln N/ac more than other choices, which seems also reflect in the worsened regret for ViOlin in Figure 7. We hypothesize that this is because, as the yield curve quickly saturates, the local curvature used for exploration in ViOlin becomes very small, causing the algorithm to underestimate uncertainty and prematurely stop exploring. Also, we notice greater variability in the estimation of parameters especially in the Logistic model as can be seen in Figure 9.
Figure 9: Parameter estimate trajectories for the well-specified setting for the Logistic model
(a) Refer to caption
(b)Refer to caption
Figure 10: Selected fertilizer rates over time for the three non-linear model based bandit algorithms for the misspecified setting. Note that the choice of fertilizer rate decreases for both UCB and ViOlin as the price per unit fertilizer increases from 0.3 to 0.7\$ lb-1N.
Misspecified setting: Additional results
For the misspecified setting, for the same combination with the true model being Mitscherlich and fitted model being Quadratic plateau, we illustrate the effect of changing the level of fertilizer price from \$0.3 to \$0.7 per unit lb N. The same hyperparameter choices are made as in the misspecified setting of Section 5.
Figure 11: Cumulative regret curves for the misspecified model corresponding to $p_{X}=0.3$ and $0.7$ respectively.
Figure 10 (a) corresponds to $p_{X}=0.3$ while Figure 10 (b) corresponds to $p_{X}=0.7$. Note that, both UCB and ViOlin capture that as the price increases the most frequently chosen arm over time becomes $x=150$ (brown dotted line) instead of $x=200$ (purple dashed dotted line). This sort of adaptive learning is a robustness check on how effective the decision-making is towards the goal of profit maximization. However, $\epsilon$-greedy failed to adapt to this change, perhaps due to a fast decaying $\epsilon_{t}$ which made it get stuck on a sub-optimal arm in the initial rounds. This is also reflected in the poor regret performance by $\epsilon$-greedy in Figure 11.