### v2

# Nonlinear Model-Based Sequential Decision-Making in Agriculture

Abstract
Agricultural decision-making faces a dual challenge: sustaining high yields to meet global food security needs while reducing the environmental impacts of fertilizer use. Nitrogen inputs are central to this tension. They are indispensable for crop growth yet also major drivers of greenhouse gas emissions, nutrient runoff, and escalating production costs. Addressing these intertwined pressures requires adaptive decision-support tools that are not only statistically principled but also economically sustainable and interpretable for practitioners.
We develop nonlinear model-based bandit algorithms as a framework for adaptive fertilizer management under uncertainty. Building on classical mechanistic yield-response models including Mitscherlich, Michaelis-Menten, quadratic plateau, and logistic functions, our approach links algorithmic exploration-exploitation strategies directly to interpretable biological processes such as maximum yield potential and nutrient efficiency. This grounding makes recommendations transparent for practitioners while supporting cost-effective and sustainable input use. Methodologically, we establish regret and sample complexity results for the well-specified nonlinear case, examine robustness under misspecification, and conduct extensive simulations reflecting realistic agricultural scenarios. The results reveal a key trade-off: nonparametric methods reduce short-term variability but introduce persistent bias, whereas model-based approaches capture biological structure and deliver more reliable long-term profit, ensuring economically sustainable fertilizer use. Our findings highlight how interpretable, uncertainty-aware algorithms can strengthen farmer livelihoods and contribute to more efficient and sustainable use of agricultural inputs.
Keywords: Sequential decision-making, agricultural statistics, nonlinear models, multi-armed bandits, resource optimization, regret
1 Introduction
Identifying the most effective crop management practice for a given site is a recurring challenge in agriculture. While field trials and expert recommendations provide useful starting points, the actual performance of management strategies such as different fertilizer rates, planting densities, or irrigation schedules can be highly variable due to differences in soil properties, previous management, and other biological or environmental factors (Hochman and Carberry,, 2011; Tilman et al.,, 2002). As a result, farmers and researchers must repeatedly decide which management actions to implement each season, often with limited prior information about their true effects. This problem exemplifies sequential decision-making under uncertainty, where the goal is to efficiently learn and optimize management choices over time in a way that supports both productivity and sustainability.
In the statistical and machine learning literature, such problems are naturally framed as multi-armed bandit problems, dating back to Robbins, (1952). They have been widely applied in areas such as precision medicine (Lu et al.,, 2021), recommendation systems (Li et al.,, 2010), and, increasingly, agriculture (Gautron et al.,, 2022). The term originates from the analogy of a gambler choosing between several slot machines (“one-armed bandits”), each with unknown payout rates. At each round, the decision-maker selects an action/arm (e.g., a specific management practice) and observes its outcome (e.g., crop yield or profit). The central challenge is to balance exploration (trying new or uncertain options to learn about their effects) with exploitation (repeating choices that have performed well so far). Bandit algorithms provide principled, data-driven strategies to manage this tradeoff, making them well-suited to resource optimization tasks in agriculture and environmental management. However, most existing bandit methods assume linear, generalized linear, or black-box (non-parametric) reward models that often fail to capture the mechanistic, nonlinear dose-response dynamics common in agronomy.
In this work, we develop and analyze a new family of bandit algorithms that directly incorporate domain-specific nonlinear response models, offering greater interpretability and sample efficiency compared to classical linear or black-box approaches. Our approach is particularly valuable in settings where repeated decisions must be made with limited data, delayed feedback, and significant economic and environmental consequences.
Nonlinear models provide parsimonious and interpretable representations of domain-specific processes. Unlike generic polynomial or linear models, their parameters often map directly to biological quantities, such as maximum yield potential or nutrient efficiency, making the models both scientifically meaningful and statistically efficient. In agriculture, such models (Miguez et al.,, 2018) capture complex responses like crop yield to fertilizer, temperature effects on growth, or pest dynamics, where the relationship follows saturating, logistic, or exponential patterns rather than a simple linear trend. Their adoption is motivated not only by statistical accuracy but also by the need for transparent recommendations in sustainable resource management
1.1 Background and related work
Multi-armed bandit algorithms (Berry and Fristedt,, 1985; Lattimore and Szepesvári,, 2018) provide a core statistical framework for sequential decision-making under uncertainty, where the key challenge is to balance exploration and exploitation. In the original bandit formulation, each arm was associated with an unknown fixed reward distribution, and repeated pulls of the same arm were assumed to yield independent and identically distributed (i.i.d.) rewards. Subsequent work extended this framework to incorporate contextual information, leading to the development of contextual bandit models (Slivkins,, 2014; Chu et al.,, 2011). In these settings, regression models such as linear (Abbasi-Yadkori et al.,, 2011), generalized linear models (GLMs) (Li et al.,, 2010; Filippi et al.,, 2010), and more flexible nonparametric approaches, including kernels, nearest neighbors, and neural networks (Yang and Zhu,, 2002; Rigollet and Zeevi,, 2010; Srinivas et al.,, 2010; Valko et al.,, 2013), enable adaptive decision-making tailored to evolving arm/covariate information.
Classical nonlinear regression models, such as Mitscherlich (Dhanoa et al.,, 2022), Michaelis-Menten (Lopez et al.,, 2000), quadratic plateau models (Bélanger et al.,, 2000), and logistic response functions (Sepaskhah et al.,, 2011), are routinely used to capture crop yield response to fertilizer and other management variables. These models offer interpretable parameters tied to underlying biological processes, which is important for transparent and actionable recommendations (De Bruin and Pedersen,, 2008; Miguez et al.,, 2018).
Machine learning approaches, including random forests, neural networks, and regression trees, have become increasingly popular for crop yield prediction and resource optimization (Shahhosseini et al.,, 2019; Khaki et al.,, 2020; Khairunniza-Bejo et al.,, 2014; Jabed and Murad,, 2024). While powerful in large-data settings, these methods are typically supervised and static, lacking the sequential and adaptive capabilities needed for decision-making under uncertainty. Concerns also remain about their interpretability and suitability for low-data environments (Dobermann et al.,, 2022).
Within agriculture, adaptive experimentation and real-time learning strategies have emerged as promising approaches for site-specific management and resource efficiency (Gautron et al.,, 2022). Recent work has explored the use of bandit algorithms (Saikai et al.,, 2020, 2018; Huang,, 2025) for adaptive plot selection and management optimization, but most existing studies focus on linear models or generic black-box methods, seldom leveraging domain knowledge of nonlinear response behavior, especially valuable for smallholders and in data-scarce scenarios (Smith,, 2018).
It is well-known that model-based bandit and reinforcement learning approaches can be far more sample-efficient than model-free methods when the underlying reward mechanisms are known (Osband and Van Roy,, 2014; Sun et al.,, 2019). However, such approaches have received limited attention in agricultural decision-making from a precision-agriculture and sustainability perspective, where established nonlinear models are often available but underutilized. Our work addresses this gap by integrating domain-driven nonlinear models directly into the bandit framework. This enables adaptive, interpretable, and sample-efficient resource management, advancing the practical and theoretical toolkit for sustainable and cost-effective fertilizer use.
Our Contributions.
To address the limitations outlined above, this paper makes the following key contributions:
  • Nonlinear model-based bandit framework: We develop a general class of multi-armed bandit algorithms that explicitly incorporate domain-specific nonlinear response functions, such as Mitscherlich and logistic models, within the exploration-exploitation loop.
  • Theoretical guarantees: We provide regret and sample complexity results under both well-specified and shape-compatible misspecified models, and discuss their implications for agricultural decision-making under uncertainty.
  • Simulation studies emulating real-world practice: Through extensive simulations inspired by realistic fertilizer optimization scenarios, we demonstrate the practical benefits of our approach over classical linear and nonparametric bandit methods, particularly in low-sample regimes.
  • Interpretability and domain relevance: By leveraging mechanistic models with interpretable parameters, our methods yield actionable recommendations for practitioners, supporting transparent, resource-efficient, and economically sustainable decision-making.
We begin by setting up the problem in Section 2. In Section 3, we lay out the proposed framework for nonlinear model-based bandits, in which we study three model-based algorithms: $\epsilon$-greedy, Upper Confidence Bound (UCB), and the ViOlin algorithm. Section 4 presents known results for parametric bandits in both well-specified and misspecified settings, and derives sample complexity bounds for the specific nonlinear models considered by the ViOlin algorithm. We then conduct extensive simulation experiments emulating real-world yield production and fertilizer optimization scenarios, with the goal of resource-efficient and cost-effective management in sample-limited settings. Finally, Section 6 concludes the paper.
2 Problem Setup: Adaptive Input Decisions with Nonlinear Models
Formally, we consider a sequential decision-making scenario with a time horizon of $T$ rounds. Each round can represent, for example, a growing season or field trial, where decisions are made at planting and outcomes observed at harvest. At each round $t=1,\dots,T$, the decision-maker (agent) chooses an action (input level) $x_{t}\in\mathcal{X}$ from a feasible set $\mathcal{X}$, such as nitrogen application rates. After choosing action $x_{t}$, the agent observes a noisy reward $y_{t}$, representing yield or profit, modeled as:
$y_{t}=f(x_{t};\beta^{*})+\epsilon_{t},$ (1)
where $f:\mathcal{X}\to\mathbb{R}$ is a known, structured nonlinear function parameterized by an unknown vector $\beta^{*}\in\mathbb{R}^{p}$, and $\epsilon_{t}$ is a mean-zero noise term. We assume $\epsilon_{t}$ is sub-Gaussian, i.e., there exists $\sigma>0$ such that for all $\lambda\in\mathbb{R}$,
$\mathbb{E}\left[e^{\lambda\epsilon_{t}}\right]\leq\exp\left(\frac{\lambda^{2}\sigma^{2}}{2}\right).$ (2)
This assumption reflects the variability of agricultural experiments driven by weather, soil heterogeneity, and biological processes, while ruling out extreme outliers that would make learning impractical.
The decision-maker seeks a policy/algorithm $\pi$ that selects actions based on historical observations to maximize cumulative expected reward. Performance can equivalently be measured through the cumulative regret, which quantifies the total loss relative to always applying the best treatment:
$R_{T}(\pi)=\sum_{t=1}^{T}\left\{f(x_{t}^{*};\beta^{*})-f(x_{t};\beta^{*})\right\},$ (3)
where $x_{t}^{*}=\arg\max_{x\in\mathcal{X}}f(x;\beta^{*})$ denotes the optimal treatment at round $t$. In agricultural terms, regret measures the cumulative yield or profit lost due to not choosing the best possible input at each season. The objective is to devise algorithms achieving sublinear regret, meaning $R_{T}(\pi)=o_{P}(T)$, so that the average regret per round diminishes over time. Such a property guarantees that the policy rapidly identifies near-optimal input levels, minimizing wasted resources and experimental costs. Other performance metrics more suitable for risk-averse decision-making have also been proposed, such as the Conditional Value-at-Risk (CVaR) of the regret, which emphasizes control of the worst-case outcomes in the tail of the regret distribution (Baudry et al.,, 2021; Wang et al.,, 2023).
To illustrate the framework, we focus on fertilizer-yield relationships and investigate four classical nonlinear models widely used in agricultural research:
  1. 1.
    Mitscherlich Model: $f(x;A,b,d)=d+A\left(1-e^{-bx}\right)$, capturing saturating yield response to fertilizer inputs.
  2. 2.
    Michaelis-Menten Model: $f(x;a,b,d)=d+\frac{ax}{b+x}$, describing nutrient uptake or growth responses with diminishing returns.
  3. 3.
    Quadratic Plateau Model:
    $f(x;a,b,c,x_{0})=\begin{cases}a+bx+cx^{2}&\text{if }x\leq x_{0}\\ a+bx_{0}+cx_{0}^{2}&\text{if }x>x_{0}\end{cases},$
    modeling responses that rise and then stabilize at a plateau, common in fertilizer trials.
  4. 4.
    Logistic Dose-Response Model: $f(x;A,B,C,d)=d+\frac{A}{1+\exp\left(-B(x-C)\right)}$, often used to capture threshold or inflection-point behavior.
These models are parsimonious, interpretable, and directly tied to biological processes. For example, in the Mitscherlich model, $A$ represents the maximum additional yield achievable with fertilizer, $b$ reflects the efficiency of nutrient uptake, and $d$ is the baseline yield. For example, a larger value of $b$ indicates that yield saturates more rapidly with increasing input, whereas a smaller $b$ corresponds to a more gradual response. Such parameters are routinely used in agronomic practice and make the resulting recommendations not only statistically grounded but also transparent and actionable.
Although we emphasize nitrogen fertilizer management as a motivating example, the framework generalizes to a wide range of agricultural and environmental problems where nonlinear dose-response relationships are well established, including irrigation efficiency, biomass accumulation, and nutrient loss in response to rainfall (Miguez et al.,, 2018). We provide detailed parameterizations and simulation examples in Section 5.
3 Methods: Algorithms for nonlinear model-based bandits
In agricultural trials, each season offers only limited opportunities for experimentation. Algorithms must therefore use past data efficiently while remaining robust to uncertainty. In this section, we describe a family of strategies for nonlinear model-based bandit problems, where the expected reward is modeled as a nonlinear function of the chosen input. We first provide a general algorithmic template, followed by detailed descriptions of three model-based strategies, namely, $\epsilon$-greedy, Upper Confidence Bound (UCB), and ViOlin, adapted to classical agronomic response models. Note that all the three algorithms have been studied in various bandit problems before (Auer and Ortner,, 2010; Arya and Yang,, 2020; Dong et al.,, 2021), here we specialize them to mechanistic nonlinear settings.
3.1 General framework
Recall, $f(x;\theta)$ denotes a known nonlinear model (e.g., Michaelis-Menten, logistic, quadratic) parameterized by $\theta$, where $x\in\mathcal{X}$ is the action (e.g., N-rate) and the expected reward is $\mathbb{E}[Y\mid x]=f(x;\theta)$. At each time $t$, the agent selects an action $x_{t}$, then receives a reward $y_{t}$.
Algorithm 1 General Nonlinear Model-based Bandit Framework
1:Input: Action set $\mathcal{X}$, nonlinear model class $f(\cdot;\theta)$, time horizon: $T$
2:Initialize: Pull each arm at least once and receive corresponding rewards until round $t_{0}$
3:for $t=t_{0}+1$ to $T$ do
4:  Estimate model parameters $\hat{\theta}_{t}$ using past data $\{(x_{s},y_{s}):s<t\}$
5:  Choose action $x_{t}$ using a strategy/policy $\pi$ utilizing $f(x;\hat{\theta}_{t})$
6:  Observe reward $y_{t}$
7:  Update dataset with $(x_{t},y_{t})$
8:end for
In the algorithm sketch (Algorithm 1), in step 3-4, at each round $t$, the decision-maker uses all previous data to update its understanding of the nonlinear reward model by estimating the unknown parameters $\theta$ that describe the model. Given past measurements of arm-reward pairs, we estimate the model parameters using nonlinear least squares, e.g., via the curve_fit function in Python’s scipy.optimize package. For a model $f(x;\theta)$ with parameters $\theta$, we solve
$\hat{\theta}=\arg\min_{\theta}\sum_{i=1}^{n}\ell(y_{i},f(x_{i};\theta)),$
where $\ell$ is a loss function (such as squared-error loss) computed using the observed $(x_{i},y_{i})$ pairs, with the goal of minimizing the loss. This practical approach is central to our implementation, allowing bandit algorithms to leverage data-driven parameter estimates and domain-specific models. Then, in step 5, the decision-maker chooses the next action by balancing the exploration-exploitation trade-off and satisfying the criterion for cumulative reward maximization. Finally, in step 6, the outcome $y_{t}$ is observed, and is added to the dataset, enabling continual refinement of parameter estimates.
Remark 1.
Unlike classical supervised learning or factorial field experiments that provide outcomes for every treatment, bandit settings reveal only the reward for the action actually chosen at each round. This partial feedback, often referred to as bandit feedback, captures the reality of agricultural experimentation where only tested management options yield observable outcomes.
Next, we describe three model-based algorithms that follow the general structure of Algorithm 1 but differ in how they balance the exploration-exploitation trade-off.
3.2 Epsilon-Greedy
The $\epsilon$-greedy algorithm (Algorithm 2) (Langford and Zhang,, 2008; Yang and Zhu,, 2002) is one of the simplest approaches yet effective approaches for balancing exploration and exploitation. At each round, the current nonlinear model is fitted using past observations, and the input that maximizes predicted profit is selected with high probability $1-\epsilon_{t}$ (exploitation). With small probability $\epsilon_{t}$, the algorithm explores by selecting a random alternative from the feasible input grid (exploration). This mechanism ensures continued learning, preventing the algorithm from becoming stuck on suboptimal choices. In agricultural contexts, this method reflects the intuitive practice of trying occasional alternative fertilizer rates even when one rate appears best.
Several variants of $\epsilon$-greedy exist; the most popular is the annealed or decaying $\epsilon$-greedy, where the exploration probability $\epsilon_{t}$ is a non-increasing sequence that tends to zero as $t\rightarrow\infty$. The intuition is that, as more data are collected and the model becomes more certain, the algorithm increasingly favors exploitation over exploration. Note that $\epsilon$ (or the sequence $(\epsilon_{t})$) is specified by the user, allowing the decision-maker to directly control the balance between exploration and exploitation; this choice directly influences the cumulative regret.
Epsilon-greedy is simple, robust, and easy to implement. It is particularly useful when the model is misspecified or when simple, interpretable algorithms are desired.
Algorithm 2 Nonlinear model-based Epsilon-Greedy
1:Input: Action set $\mathcal{X}$, Exploration parameter $\epsilon>0$, non-linear model class $f(\cdot;\theta)$
2:Initialize: Random pull arms and receive corresponding rewards until round $t_{0}$
3:for $t=t_{0}+1$ to $T$ do
4:  Estimate $\hat{\theta}_{t}$ from previous data: $\{x_{1},y_{1},\ldots,x_{t-1},y_{t-1}\}$
5:  Select action $x_{t}$ according to:
$x_{t}=\begin{cases}\text{sample uniformly from }\mathcal{X}&\text{with probability }\epsilon\\ \arg\max_{x\in\mathcal{X}}f(x;\hat{\theta}_{t})&\text{with probability }1-\epsilon\end{cases}$
6:  Observe $y_{t}$, update data
7:end for
If $\mathcal{X}$ is a finite set of $K$ elements, “sample uniformly from $\mathcal{X}$” means selecting an element at random with equal probability, so that each arm is chosen with probability $1/K$. In the $\epsilon$-greedy algorithm, during the exploration step, each arm is selected with probability $\epsilon/K$. If $\mathcal{X}$ is a continuous interval or region, this means sampling from the continuous uniform distribution over $\mathcal{X}$, i.e., $x_{t}\sim\mathrm{Unif}(\mathcal{X})$.
3.3 Nonlinear model-based UCB
Overview:
The UCB (Upper Confidence Bound) algorithm (Auer and Ortner,, 2010; Chu et al.,, 2011; Zhou et al.,, 2020) embodies the principle of “optimism in the face of uncertainty” by selecting actions that maximize an upper confidence bound on the expected reward. Here, the nonlinear model is repeatedly fitted, and for each input level the algorithm constructs an upper confidence bound on the predicted reward. The next action is chosen to maximize this upper bound. This mechanism naturally prioritizes inputs that are either promising in terms of mean yield or remain highly uncertain, thereby encouraging exploration in scientifically justified directions. Applied to fertilizer management, UCB corresponds to testing input levels where either predicted yield is high or parameter uncertainty remains large.
In model-based nonlinear regression, the uncertainty $\mathrm{Unc}_{t}(x)$ is often based on the variability of the estimated reward due to finite data, as quantified via the standard error of the model prediction. Specifically, the uncertainty term is given by
$\mathrm{Unc}_{t}(x)=\sqrt{\nabla_{\theta}f(x;\hat{\theta}_{t})^{\top}\,\widehat{\mathrm{Cov}}(\hat{\theta}_{t})\,\nabla_{\theta}f(x;\hat{\theta}_{t})},$ (4)
where $\nabla_{\theta}f(x;\hat{\theta}_{t})$ denotes the gradient of the reward model with respect to the parameters, evaluated at the current estimate $\hat{\theta}_{t}$, and $\widehat{\mathrm{Cov}}(\hat{\theta}_{t})$ is the estimated covariance matrix of $\hat{\theta}_{t}$.
The UCB action selection rule then takes the form
$x_{t}=\arg\max_{x\in\mathcal{X}}\,f(x;\hat{\theta}_{t})+\alpha\cdot\mathrm{Unc}_{t}(x)$ (5)
where $\alpha>0$ is a user-specified confidence parameter that controls the degree of exploration.
UCB algorithms provide strong theoretical guarantees for balancing exploration and exploitation, and are particularly effective when reliable, model-based estimates of prediction uncertainty are available (see, e.g., (Zhou et al.,, 2020)). It is particularly advantageous in settings where data collection is expensive or risky, as it prioritizes actions that could yield high rewards or that have not been thoroughly explored. Model-based UCB (Algorithm 6) thus combines statistical rigor with the agronomic interpretability of nonlinear functions, making it a strong candidate for adaptive on-farm experimentation.
Algorithm 3 Nonlinear model-based UCB
1:Input: Action set $\mathcal{X}$, confidence parameter $\alpha>0$
2:for $t=1$ to $T$ do
3:  Estimate $\hat{\theta}_{t}$ from the previous data
4:  for each $x\in\mathcal{X}$ do
5:   Compute predicted reward $f(x,z_{t};\hat{\theta}_{t})$ and uncertainty $\mathrm{Unc}_{t}(x)$ as defined in (4)
6:   Calculate $\text{UCB}_{t}(x)=f(x,z_{t};\hat{\theta}_{t})+\alpha\cdot\mathrm{Unc}_{t}(x)$
7:  end for
8:  Choose $x_{t}=\arg\max_{x\in\mathcal{X}}\text{UCB}_{t}(x)$
9:  Observe $y_{t}$, update data
10:end for
3.4 ViOlin (Virtual Ascent with Online Model Learner)
The ViOlin algorithm (Algorithm 4(Dong et al.,, 2021) is a model-based bandit strategy designed for efficient learning in nonlinear settings. Unlike $\epsilon$-greedy, which explores actions uniformly at random, or UCB, which prioritizes actions with high uncertainty, ViOlin is a greedy method that leverages both the current model estimate and local geometric information about the reward surface (i.e., gradient and curvature). At each round, the algorithm chooses the action predicted to be the best for profit maximization by the current estimated model, then it updates the model using not only the observed rewards, but also estimates of the local gradient and curvature of the reward function at the selected action. This curvature-matching approach yields strong sample efficiency guarantees, especially in complex, high-dimensional, or highly nonlinear problems, by ensuring the model learns both the value and the local shape of the reward function. While originally, Dong et al., (2021) present the ViOlin algorithm for the non-noisy setting, we illustrate its performance under model noise as well.
Algorithm 4 ViOlin: Virtual Ascent with Online Model Learner (Dong et al.,, 2021)
1:Input: Model class $\mathcal{F}$ parameterized by $\theta$, action set $\mathcal{X}$, initial model $\hat{\theta}_{1}$, online learner $\mathcal{O}$, total rounds $T$
2:for $t=1$ to $T$ do
3:  Model Update: Update model parameters $\hat{\theta}_{t}$ using all data observed so far via the online learner $\mathcal{O}$
4:  Action Selection: Choose
$x_{t}=\arg\max_{x\in\mathcal{X}}f(x;\hat{\theta}_{t})$
5:  Play and Observe: Play action $x_{t}$, observe the true reward $r_{t}=f(\theta^{*},x_{t})+\epsilon_{t}$
6:  Curvature Estimation: Estimate the gradient and Hessian of the real reward function at $x_{t}$ using random perturbations or finite differences:
$\widehat{\nabla}_{x}f(\theta^{*},x_{t}),\quad\widehat{\nabla}^{2}_{x}f(\theta^{*},x_{t})$
7:  Curvature-Matching Model Update: Update $\mathcal{O}$ by minimizing
$\ell_{t}(\theta)=\;$ $(f(x_{t};\theta)-y_{t})^{2}+\alpha_{1}\left\|\nabla_{x}f(x_{t};\theta)-\widehat{\nabla}_{x}f^{*}(x_{t})\right\|^{2}$
$+\alpha_{2}\left\|\nabla^{2}_{x}f(x_{t};\theta)-\widehat{\nabla}^{2}_{x}f^{*}(x_{t})\right\|^{2}$
8:end for
ViOlin is particularly attractive in data-limited settings (e.g., smallholder or resource-constrained agricultural experiments), where minimizing the number of field trials is crucial. ViOlin’s greedy exploration is guided by the model fit and its local geometry, without explicit randomized or uncertainty-driven action selection. This enables fast convergence to (local) optima when the model class matches the true reward structure and curvature information is reliable. In low-dimensional agronomic models with smooth response surfaces, ViOlin performs on par with simpler strategies such as $\epsilon$-greedy and UCB. Its strength becomes pronounced as complexity grows: in multi-modal settings or higher-dimensional decision spaces, ViOlin’s sample-efficiency guarantees provide a clear advantage. This makes ViOlin well suited for adaptive experimentation in limited data regimes, such as on-farm field trials or precision agriculture for smallholder contexts, where each data point is costly to obtain.
Summary of algorithm differences:
$\epsilon$-greedy explores uniformly at random, UCB targets actions with high uncertainty (optimism-based exploration), and ViOlin is greedy but uses strong model-based curvature matching for sample-efficient learning in complex settings.
To illustrate how these algorithms operate and compare in practice, we include a simple, step-by-step example using the Mitscherlich model over three or four rounds in Section 3 of the Appendix. This concrete illustration should help readers, especially those less familiar with bandit algorithms, understand the exploration-exploitation tradeoff.
While the primary emphasis of this work is not on theoretical development, we present and discuss several well-known theoretical results and highlight their implications for the nonlinear model-based bandit framework.
4 Theoretical guarantees for nonlinear bandit algorithms
In this section, we characterize the regret and sample-complexity guarantees of sequential algorithms under increasingly flexible reward models, beginning with classical linear formulations, extending to nonparametric function classes, then focusing on nonlinear parametric (mechanistic) models as used in agronomy, and finally accounting for model misspecification. While the general theory applies broadly, our emphasis is on sample efficiency and robustness in low-data agricultural regimes, where each field trial is costly and decision errors translate to both economic and environmental consequences.
Linear and GLM bandits.
For linear bandits with $d$-dimensional parameterization, the minimax expected cumulative regret is known to satisfy
$\text{E}[R_{T}]=O\left(d\sqrt{T}\log T\right),$
where $T$ is the time horizon (Abbasi-Yadkori et al.,, 2011; Dani et al.,, 2008). Generalized linear bandit models achieve similar rates up to logarithmic factors (Filippi et al.,, 2010; Bastani et al.,, 2020). These results provide a baseline: when yield or profit responds linearly to inputs, bandit algorithms can achieve fast learning rates. However, such linear approximations are rarely biologically realistic in crop response.
Nonparametric bandits.
For more general nonparametric function classes, such as Lipschitz or RKHS (kernelized) reward functions, the minimax regret scales polynomially with the dimension $d$ of the action/context space $\mathcal{X}$  (Slivkins,, 2014; Srinivas et al.,, 2012):
$R_{T}=\widetilde{O}\left(T^{\frac{d+1}{d+2}}\right).$
While attractive for flexibility, these guarantees deteriorate with dimension and are impractical when only a small number of seasons or trials are available, which is a common reality in agriculture.
NeuralUCB and expressive function classes.
Recent work has extended bandit algorithms to highly expressive nonlinear function classes via neural networks. For instance, NeuralUCB (Zhou et al.,, 2020) achieves regret rates of order
$R_{T}=\widetilde{O}\!\left(\tilde{d}(\lambda)\sqrt{T}\right),$
where $\tilde{d}(\lambda)$ is the effective dimension of the neural tangent kernel associated with the underlying network. For example, for a two-layer ReLU neural network with input dimension $d$ and width $m$, it holds that $\tilde{d}(\lambda)\leq d\,\mathrm{poly}(m,\log(T/\lambda))$, so that the regret rate can be written as
$R_{T}=\widetilde{O}\left(d\,\mathrm{poly}(m)\sqrt{T}\right).$
Here, $\mathrm{poly}(m,\cdot)$ denotes a polynomial function of the network width and the indicated arguments. Thus, the regret bound increases with both the expressiveness of the neural network (through $m$) and the ambient input dimension $d$, but retains the $\sqrt{T}$ scaling characteristic of parametric (linear) bandit models. However, these improvements come with heavy computational costs, the need for large data, and limited interpretability.
We summarize the regret bounds for the four classes of bandit problems in Table 1. While such results highlight the theoretical reach of modern methods, they are poorly suited to agricultural decision-making, where data are scarce, feedback is seasonal, and recommendations must be transparent to practitioners. Our focus, therefore, is on mechanistic yield-response models such as Mitscherlich or Michaelis-Menten. These models are parsimonious, interpretable, and biologically grounded, yet flexible enough to capture crop dose-response behavior. This specialization ensures theoretical guarantees translate into actionable insights, directly relevant for resource-efficient agricultural experimentation.
Table 1: Summary of Regret Rates for Bandits under Different Model Classes
Model Class Regret Bound Reference
Linear Bandit $O(d\sqrt{T}\log T)$ Abbasi-Yadkori et al., (2011)
GLM Bandit $\widetilde{O}(d\sqrt{T})$ Filippi et al., (2010)
Nonparametric (Lipschitz, RKHS) $\widetilde{O}(T^{\frac{d+1}{d+2}})$ Slivkins, (2014)
NeuralUCB (2-layer NN) $\widetilde{O}(d\,\mathrm{poly}(m)\sqrt{T})$ Zhou et al., (2020)
4.1 Model-based nonlinear bandits.
Building on these insights, we now focus on model-based nonlinear bandits, a regime particularly relevant in agronomy, where biological processes often admit interpretable, low-dimensional parameterizations. Since one of our motivations for this work is to provide decision-making algorithms in sample-limited scenarios, we focus on characterizing sample complexity more than providing regret guarantees in this section. Recent work by Dong et al., (2021) has shown a breakthrough for model-based nonlinear bandits: for any class of reward functions with bounded sequential Rademacher complexity, it is possible to find an $\epsilon$-approximate local maximum with sample complexity polynomial in the complexity of the model class, independent of the action dimension. This suggests that, unlike classical methods whose sample complexity or regret is exponential in dimension, model-based methods can be vastly more efficient when the reward function admits a suitable low-complexity parameterization.
Sample Complexity and Local Regret for Model-based Nonlinear Bandits.
In the model-based bandit setting, the functional form of the reward is known, but its parameters are not. The goal is therefore to efficiently identify an approximate local maximum of this parametric reward function, rather than a global maximum which is often computationally and statistically intractable. Here, sample complexity refers to the minimum number of experimental rounds required to guarantee, with high probability, that the recommended action achieves performance within a specified tolerance of a local optimum. This guarantee is formalized through the notion of local regret, which measures the suboptimality of the chosen action relative to the best locally optimal action. In practice, for well-behaved (e.g., unimodal) functions such as those common in agronomic yield response, a local maximum often coincides with the global maximum.
In particular, a point $x$ is said to be an $(\epsilon_{g},\epsilon_{h})$-approximate local maximum if its gradient is at most $\epsilon_{g}$ (i.e., $\|\nabla f(x)\|_{2}\leq\epsilon_{g}$) and the Hessian’s largest eigenvalue is at most $\epsilon_{h}$ (i.e., $f$ is sufficiently concave around $x$). The sample complexity quantifies how quickly an algorithm can find such a point.
In this work, we focus on reward functions (e.g., quadratic-plateau, Michaelis-Menten, Mitscherlich, and logistic) that satisfy the regularity conditions (such as bounded gradients and Hessians) assumed in recent theoretical results (Dong et al.,, 2021). These properties ensure that our algorithms are well-behaved: the models are smooth enough for efficient learning, and the theoretical sample complexity guarantees apply. As a result, we can meaningfully compare algorithms in terms of how many rounds are needed to reach near-optimal fertilizer recommendations with high confidence. We first define the notion of sequential Rademacher complexity and then state Theorem 1.1 from Dong et al., (2021) in terms of sequential complexity, and then we tailor the result to our specific non-linear model classes.
Intuitively, sequential Rademacher complexity measures how hard it is for a learning algorithm to reliably make good decisions when facing an environment that can adapt to the algorithm’s past actions. More formally, the sequential Rademacher complexity (Rakhlin et al., 2015b, ) is defined as follows.
Definition 1 (Sequential Rademacher Complexity).
Let $\mathcal{F}$ be a class of real-valued functions defined on $\mathcal{X}$. The sequential Rademacher complexity of $\mathcal{F}$ over $T$ rounds is defined as
$\mathfrak{R}_{T}^{\mathrm{seq}}(\mathcal{F})=\sup_{x_{1},\ldots,x_{T}}\mathbb{E}_{\epsilon}\left[\sup_{f\in\mathcal{F}}\frac{1}{T}\sum_{t=1}^{T}\epsilon_{t}f\left(x_{t}(\epsilon_{1},\ldots,\epsilon_{t-1})\right)\right]$
where the supremum is taken over all sequences of functions $x_{t}:\{-1,+1\}^{t-1}\to\mathcal{X}$, and $\epsilon_{1},\dots,\epsilon_{T}$ are independent Rademacher random variables, i.e., $\mathbb{P}(\epsilon_{t}=1)=\mathbb{P}(\epsilon_{t}=-1)=\frac{1}{2}$ for each $t$.
Now, we present the result by Dong et al., (2021) that determines the sample complexity rate for ViOlin.
Theorem 1 (Sample Complexity for Model-based Nonlinear Bandits (Dong et al.,, 2021, Theorem 1.1)).
Suppose the sequential Rademacher complexity of the loss function class $\mathfrak{R}_{T}^{\mathrm{seq}}(\mathcal{F})$ induced by the reward function class $\{f(\theta,\cdot):\theta\in\Theta\}$ is bounded by $\sqrt{R(\Theta)T\,\mathrm{polylog}(T)}$ for some complexity parameter $R(\Theta)$. Then, there exists an algorithm (ViOlin)) (Algorithm  4) that finds an $\epsilon$-approximate local maximum with $\widetilde{O}(R(\Theta)\epsilon^{-8})$ samples, independent of the dimension of the action space.
Note that most of the agronomic models assume bounded non-linear functions, such as ones that exhibit plateauing behavior after a certain threshold. Therefore, it is of interest to quantify the sequential Rademacher complexity for these class of functions. For parametric classes with bounded functions and inputs (such as those considered in this paper), we prove a theoretical bound on the sequential Rademacher complexity of bounded functions and show that indeed for this class of functions the sequential Rademacher complexity has an upper bound as required in Theorem 1.
Theorem 2 (Sequential Rademacher Complexity for Bounded Functions).
Let $\mathcal{F}\subseteq[-B,B]^{\mathcal{D}}$ be a class of functions uniformly bounded by $B>0$. Then the sequential Rademacher complexity of $\mathcal{F}$ satisfies
$\mathfrak{R}_{T}^{\mathrm{seq}}(\mathcal{F}):=\sup_{\boldsymbol{x}}\mathbb{E}_{\boldsymbol{\epsilon}}\left[\sup_{f\in\mathcal{F}}\frac{1}{T}\sum_{t=1}^{T}\epsilon_{t}f(x_{t}(\epsilon_{1},\ldots,\epsilon_{t-1}))\right]\leq C\cdot B\cdot\sqrt{\frac{\log T}{T}},$
for some universal constant $C>0$.
We defer the proof of Theorem 2 to the Appendix (Section A) for brevity and discuss its implications from an application viewpoint. Our theoretical result complement Dong et al., (2021)’s general framework by concretely instantiating the sample complexity for these practically important nonlinear classes, including quadratic plateau (threshold), Michaelis-Menten, Mitscherlich, and logistic models, the sequential Rademacher complexity $\mathfrak{R}_{T}^{\mathrm{seq}}(\mathcal{F})$ can be bounded as $\widetilde{O}(1/\sqrt{T})$. Therefore, the sample complexity for identifying an $\epsilon$-optimal arm (in the sense of a local maximum) in these bandit settings scales as $\widetilde{O}(1/\epsilon^{8})$, with constants depending on the parameter and input bounds. This bound, as established in Dong et al., (2021), is general and not specialized to the structure of agronomic models in terms of its polynomial dependence on $\epsilon$. Deriving sharper rates for the structured nonlinear models considered here is an important direction for future work.
Corollary 1 (Sample Complexity for online learning of Bounded Non-linear Reward classes).
Using the bound in Theorem 2 in Theorem 1, an $\epsilon$-optimal solution (arm) can be identified with sample complexity (number of time steps), $\widetilde{O}(1/\epsilon^{8})$ for the ViOlin algorithm of (Dong et al.,, 2021), independent of the action (arm) space dimension.
Interpretation.
This shows that for commonly used nonlinear agronomic yield response models, bandit algorithms can efficiently identify near-optimal fertilizer rates with a number of field trials that scales polynomially with the desired accuracy and avoids the curse of dimensionality in the action space. Our results instantiate Dong et al., (2021)’s general theory for these specific models, providing practical sample complexity bounds for adaptive experimentation in agriculture. In Theorem 1, the parameter $\epsilon>0$ represents the maximum allowable gap between the algorithm’s selected action (e.g., a nitrogen rate recommendation) and the optimal (locally maximizing) action in terms of expected reward. That is, with high probability, the bandit algorithm identifies an action $x$ such that
$f^{*}(x^{*})-f^{*}(x)\leq\epsilon,$
where $f^{*}$ is the true (unknown) reward or profit function, $x^{*}$ is a (local) maximizer of $f^{*}$, and $x$ is the action recommended by the algorithm.
The sample complexity required to achieve an $\epsilon$-optimal recommendation scales polynomially with $1/\epsilon$ (specifically, as $O(1/\epsilon^{8})$ in Theorem 1), so tighter tolerances (smaller $\epsilon$) require proportionally more field trials. In practice, $\epsilon$ should be set according to what constitutes an economically meaningful margin for decision-making.
Dong et al., (2021) provide broad sample complexity and regret guarantees for nonlinear bandit problems, including highly expressive classes such as 2-layer neural networks. By contrast, our agricultural decision models are low-dimensional and mechanistic (e.g., quadratic plateau), which makes them both interpretable and practically relevant. In this structured setting, algorithms such as UCB and $\epsilon$-greedy are not only straightforward to implement but also expected to admit stronger theoretical guarantees than the general results in Dong et al., (2021). We adopt their bounds as a reference baseline, while highlighting the opportunity for future work to establish sharper, model-specific results. This positions our empirical study on firm theoretical footing while opening avenues for more refined analysis.
Remark 2.
The proof techniques used here such as bounding regret via sequential Rademacher complexity and the resulting sample-complexity corollary for ViOlin are direct instantiations of general tools from the bandit literature. What is distinctive in this result is their adaptation to mechanistic nonlinear models (e.g., Mitscherlich, Michaelis-Menten), which integrates domain knowledge with statistical guarantees. This integration ensures that theoretical results are not only mathematically sound but also directly interpretable and actionable in agronomic applications.
Under Model Misspecification
In many applications, including agricultural yield optimization, the true reward function may not be perfectly captured by any member of a chosen parametric model class (such as quadratic plateau or Michaelis-Menten). This is known as model misspecification, and can have important implications for the performance and guarantees of bandit algorithms.
Most bandit algorithms rely on the assumption of a well-specified model (or “realizability”), that is, the existence of a function $f^{*}\in\mathcal{F}$ such that the mean reward is $\mu(x)=f^{*}(x)$ for all actions $x$ (Chu et al.,, 2011; Abbasi-Yadkori et al.,, 2011; Agarwal et al.,, 2012; Foster et al.,, 2018). However, in practice, exact realizability rarely holds, and it is important to consider the effect of model misspecification.
Recent work (see, e.g., (Greenewald et al.,, 2021; Lattimore et al.,, 2020; Ghosh et al.,, 2017)) has studied bandits under various notions of misspecification. A common formulation is the uniform $\varepsilon$-misspecification setting, where
$\inf_{f\in\mathcal{F}}\sup_{x\in\mathcal{X}}\left|\mu(x)-f(x)\right|\leq\varepsilon,$ (6)
for some $\varepsilon>0$.
For linear bandits, Lattimore et al., (2020) show that the cumulative regret must satisfy
$R_{T}\gtrsim d\sqrt{T}+\varepsilon\sqrt{d}T,$
where $d$ is the model dimension and $T$ is the time horizon. The first term corresponds to the minimax regret under well-specified models, while the second is an unavoidable “price of misspecification”. In our setting, the models (quadratic plateau, logistic, etc.) may be misspecified relative to the true reward. As such, the regret of our bandit algorithms can be interpreted as
$R_{T}=O(\text{model-based regret})+T\cdot\varepsilon_{T}(S),$
where the first term is the regret under realizability, and the second term is proportional to the sequence-specific misspecification error. This highlights the need for robust algorithms and motivates future work on model selection or hybrid model-based/nonparametric approaches.
5 Simulation study & Results: Emulating real-world agricultural decision-making
5.1 From yield to profit: Economic optimization of fertilizer
In agricultural decision-making, maximizing yield alone does not necessarily lead to optimal outcomes. Most agronomic yield functions exhibit diminishing returns to fertilizer input, i.e., yield increases with fertilizer up to a point and then plateaus. Since fertilizer is a costly input, applying more of it beyond the agronomically efficient range leads to higher expenses without commensurate gains in yield. As a result, the economically optimal fertilizer dose is not the one that produces the maximum yield, but rather the one that maximizes profit, the difference between revenue and input cost.
To formalize this, let $p_{y}$ denote the price per unit yield (e.g., \$/bushel), and $p_{x}$ the price per unit of fertilizer (e.g., \$/lb N). Suppose the yield response is described by a parametric function $f(x;\theta)$, where $x$ is the fertilizer application level and $\theta$ denotes model parameters capturing biological or agronomic characteristics. The corresponding economic profit is then given by
$\Pi(x)=p_{y}\cdot f(x;\theta)-p_{x}\cdot x.$ (7)
The economically optimal fertilizer level $x^{*}$ maximizes this profit:
$x^{*}=\arg\max_{x\in\mathcal{X}}\,\Pi(x)=\arg\max_{x\in\mathcal{X}}\,\left[p_{y}f(x;\theta)-p_{x}x\right],$ (8)
where $\mathcal{X}$ is the feasible set of fertilizer rates (e.g., a grid from 0 to 250 lb N/ac). Depending on the form of $f$, this maximization may admit a closed-form solution or may require numerical optimization. These generic routines can accommodate any parametric or nonparametric yield response function $f(x;\theta)$, enabling interpretable economic optimization across a broad range of agronomic scenarios. For the nonlinear yield functions studied in our simulations, closed-form expressions for $x^{*}$ can be derived and are summarized in Table 2. These expressions highlight how the optimal fertilizer rate depends jointly on agronomic parameters $\theta$ and economic parameters $p_{y},p_{x}$, all of which must be estimated from data.
We now turn to the problem of sequentially optimizing fertilizer decisions using model-based bandit algorithms that account for uncertainty and limited data. At each round, the algorithm fits a nonlinear yield response model to the observed data and selects the fertilizer rate that maximizes the estimated economic profit, optionally incorporating an uncertainty-based exploration bonus. We investigate three nonlinear model-based strategies:
  1. 1.
    Model-based $\epsilon$-greedy algorithm (Algorithm 5): This method selects a random fertilizer rate with probability $\epsilon_{t}=t^{-a}$, with $a$ chosen from $\{0.5,1,1.5\}$, for exploration. With the remaining $(1-\epsilon_{t})$ probability, it chooses the rate closest to the current profit-maximizing estimate $x^{*}$, computed using the fitted model and closed-form expressions where available.
  2. 2.
    Model-based UCB algorithm (Algorithm 6): This approach augments the estimated profit at each arm with a model-based uncertainty term, derived from the parameter covariance of the fitted nonlinear model, and selects the rate with the highest upper confidence bound.
  3. 3.
    ViOlin) algorithm (Algorithm 4): A curvature-aware strategy that exploits local second-order structure in the profit function to accelerate convergence, particularly under smooth nonlinearity.
Although these methods differ in their exploration strategies, they all share the same structure: fitting a nonlinear model to guide adaptive decision-making under uncertainty. Accurate initialization of nonlinear model parameters is crucial to the success of these algorithms, especially in small-sample settings. To benchmark the performance of nonlinear model-based strategies, we compare them against two widely used baselines: LinUCB (Chu et al.,, 2011) and $k$-Nearest Neighbor-UCB (kNN-UCB) (Reeve et al.,, 2018).
  1. 4.
    LinUCB algorithm (Algorithm 7): A linear bandit method that models yield as a linear function of fertilizer rate and selects actions using upper confidence bounds derived from online regression.
  2. 5.
    kNN-UCB algorithm (Algorithm 8): A nonparametric method that estimates yield using the $k$ nearest neighbors of each candidate arm and quantifies uncertainty via local sample variance.
Further descriptions of these algorithms can be found in the Appendix. These baselines span the spectrum from parametric simplicity to nonparametric flexibility and help highlight the advantages of domain-informed nonlinear models in sample-constrained settings.
Table 2: Closed-form expressions for profit-maximizing fertilizer doses under different nonlinear yield models. These highlight how optimal input use depends jointly on agronomic and economic parameters, providing interpretable decision rules.
Model Yield Function $Y(x)$ Profit-maximizing $x^{*}$
Mitscherlich $A(1-e^{-bx})$ $\frac{1}{b}\ln\left(\frac{p_{x}}{p_{y}Ab}\right)$
Quadratic (threshold) $\begin{array}[]{l}a+bx+cx^{2},\ x\leq x_{0}\\ a+bx_{0}+cx_{0}^{2},\ x>x_{0}\end{array}$ $\min\left\{x_{0},\ \max\left\{0,\ \frac{1}{2c}\left(\frac{p_{x}}{p_{y}}-b\right)\right\}\right\}$
Michaelis-Menten $\dfrac{ax}{b+x}$ $\sqrt{\dfrac{abp_{y}}{p_{x}}}-b$
Logistic $\dfrac{A}{1+e^{-B(x-C)}}$ $\begin{array}[]{l}u^{*}=\dfrac{\gamma-2-\sqrt{(\gamma-2)^{2}-4}}{2},\\ \gamma=\dfrac{Bp_{y}A}{p_{x}},\quad x^{*}=C-\dfrac{1}{B}\ln u^{*}\end{array}$
Algorithm 5 Model-Based $\epsilon$-Greedy for Economic Profit Maximization
0: Fertilizer levels $\mathcal{X}$, time horizon $T$, prices $p_{y},p_{x}$, exploration rate $\epsilon_{t}$
1: Initialize dataset $\mathcal{D}\leftarrow\emptyset$
2:for $t=1$ to $T$ do
3:  if $u\sim\text{Uniform}(0,1)<\epsilon_{t}$ or $|\mathcal{D}|<n_{0}$ then
4:   Select fertilizer level $x_{t}\sim\text{Uniform}(\mathcal{X})$ // Explore
5:  else
[…]
7:   Compute $x^{*}=\arg\max_{x\in\mathcal{X}}\,\left[p_{y}f(x;\hat{\theta})-p_{x}x\right]$
8:   Select $x_{t}\in\mathcal{X}$ closest to $x^{*}$
9:  end if
10:  Apply $x_{t}$, observe yield $Y_{t}$
11:  Compute profit: $\Pi_{t}=p_{y}\cdot Y_{t}-p_{x}\cdot x_{t}$
12:  Add $(x_{t},Y_{t})$ to dataset $\mathcal{D}$
13:end for
Algorithm 6 Model-Based UCB for Economic Profit Maximization
0: Fertilizer levels $\mathcal{X}$, time horizon $T$, prices $p_{y},p_{x}$, UCB constant $\alpha$
1: Initialize dataset $\mathcal{D}\leftarrow\emptyset$
2:for $t=1$ to $T$ do
[…]
5:  for each $x\in\mathcal{X}$ do
6:   Predict profit: $\hat{\Pi}(x)=p_{y}\cdot f(x;\hat{\theta})-p_{x}\cdot x$
7:   Compute model-based uncertainty:
$\mathrm{Unc}(x)=p_{y}\cdot\sqrt{\nabla_{\theta}f(x;\hat{\theta})^{\top}\widehat{\mathrm{Cov}}(\hat{\theta})\nabla_{\theta}f(x;\hat{\theta})}$
8:   Compute UCB score: $UCB(x)=\hat{\Pi}(x)+\alpha\cdot\mathrm{Unc}(x)$
9:  end for
10:  Select $x_{t}=\arg\max_{x\in\mathcal{X}}UCB(x)$
11:  Apply $x_{t}$, observe yield $Y_{t}$
12:  Compute profit: $\Pi_{t}=p_{y}\cdot Y_{t}-p_{x}\cdot x_{t}$
13:  Update $\mathcal{D}\leftarrow\mathcal{D}\cup\{(x_{t},Y_{t})\}$
14:end for
Algorithm 7 LinUCB for Economic Profit Maximization
0: Fertilizer levels $\mathcal{X}$, time horizon $T$, prices $p_{y},p_{x}$, UCB parameter $\alpha$
1: Initialize dataset $\mathcal{D}\leftarrow\emptyset$
2: Define feature mapping $\varphi(x)=[1,\,x]^{\top}$
3:for $t=1$ to $T$ do
4:  if not enough data then
5:   Choose $x_{t}$ uniformly at random from $\mathcal{X}$
6:  else
7:   Fit linear model to data: $Y_{i}\approx\varphi(x_{i})^{\top}\hat{\beta}$
8:   For each $x\in\mathcal{X}$:
  • Predict yield: $\widehat{Y}(x)=\varphi(x)^{\top}\hat{\beta}$
  • Compute profit: $\hat{\Pi}(x)=p_{y}\cdot\widehat{Y}(x)-p_{x}\cdot x$
  • Estimate variance: $s^{2}(x)=\varphi(x)^{\top}(\mathbf{V}^{-1})\varphi(x)$, where $\mathbf{V}$ is the feature covariance matrix
  • Compute UCB score: $UCB(x)=\hat{\Pi}(x)+\alpha\cdot s(x)$
9:   Select $x_{t}=\arg\max_{x\in\mathcal{X}}UCB(x)$
10:  end if
11:  Apply $x_{t}$, observe $Y_{t}$
12:  Compute profit: $\Pi_{t}=p_{y}\cdot Y_{t}-p_{x}\cdot x_{t}$
13:  Append $(x_{t},Y_{t})$ to $\mathcal{D}$
14:end for
Algorithm 8 kNN-UCB for Economic Profit Maximization
0: Fertilizer levels $\mathcal{X}$, time horizon $T$, prices $p_{y},p_{x}$, UCB parameter $\alpha$, no. of neighbors $k$
1: Initialize dataset $\mathcal{D}\leftarrow\emptyset$
2:for $t=1$ to $T$ do
3:  if not enough data then
4:   Choose $x_{t}$ uniformly at random from $\mathcal{X}$
5:  else
6:   for each $x\in\mathcal{X}$ do
7:    Find $k$ nearest previously tried fertilizer rates to $x$ in $\mathcal{D}$
8:    Compute average yield: $\widehat{Y}_{k}(x)=\frac{1}{k}\sum_{j=1}^{k}Y_{j}$
9:    Compute sample standard deviation: $s_{k}(x)$ of yields among neighbors
10:    Compute profit: $\hat{\Pi}_{k}(x)=p_{y}\cdot\widehat{Y}_{k}(x)-p_{x}\cdot x$
11:    Compute UCB score: $UCB(x)=\hat{\Pi}_{k}(x)+\alpha\cdot\frac{s_{k}(x)}{\sqrt{k}}$
12:   end for
13:   Select $x_{t}=\arg\max_{x\in\mathcal{X}}UCB(x)$
14:  end if
15:  Apply $x_{t}$, observe $Y_{t}$
16:  Compute profit: $\Pi_{t}=p_{y}\cdot Y_{t}-p_{x}\cdot x_{t}$
17:  Append $(x_{t},Y_{t})$ to $\mathcal{D}$
18:end for
In the next section, we present simulation studies under both well-specified and misspecified conditions to compare the performance of these five bandit algorithms in terms of profit and cumulative regret.
5.2 Well-specified setting
Simulation setup.
In the well-specified setting, data were generated according to a quadratic plateau model of the form:
$Y(x)=\begin{cases}a+bx+cx^{2},&\text{if }x\leq x_{0}\\ a+bx_{0}+cx_{0}^{2},&\text{if }x>x_{0}\end{cases},$
where $Y(x)$ denotes crop yield (bu/ac) at fertilizer rate $x$ (lb N/ac), and $a=80$, $b=1.2$, c=0.003c=-0.003, $x_{0}=180$ were chosen to reflect realistic agronomic response curves. The profit at each round is computed as $\Pi(x)=p_{y}Y(x)-p_{x}x$, with grain price $p_{y}=\$5.00$ per bushel and fertilizer price $p_{x}\in\{0.3,0.5,0.7\}$ \$/lb N to mimic realistic prices per unit nitrogen fertilizer (such as Urea) in the Midwest US. Fertilizer rates are restricted to the discrete grid $\mathcal{X}=\{0,50,100,\ldots,250\}$ lb N/ac, and yields are corrupted by i.i.d. Gaussian noise with standard deviation $\sigma=0.5$.
We evaluated five algorithms: model-based $\epsilon$-greedy, nonlinear UCB, ViOlin (curvature-matched nonlinear bandit), LinUCB (linear model-based UCB), and kNN-UCB (nonparametric UCB using $k=3$ nearest neighbors). Hyperparameters for each of the algorithms are chosen based on theoretical guidelines and grid searches over multiple runs. We choose $\epsilon_{t}=t^{-1.5}$ for $\epsilon$-greedy in this setting, $\alpha=1$ in UCB, linUCB and kNN-UCB for uniformity and fairness in comparison, $\alpha_{1}=2.0,\alpha_{2}=640$ for the ViOlin algorithm based on guidelines in Dong et al., (2021). For each algorithm and parameter configuration, we performed $10$ independent simulation replicates of $T=30$ rounds. Each algorithm fits its specified model to the observed data and selects actions to maximize estimated profit. The regret at each round is defined as the difference between the average profit for the optimal arm (computed using knowledge of the true reward function) and the average profit realized by the arm selected by the algorithm, that is,
$\mathrm{Regret}_{t}=\mathbb{E}\left[\Pi(x^{*};\theta^{*})-\Pi(x_{t};\theta^{*})\right],$
where $x^{*}$ denotes the arm maximizing the true profit function, $\theta^{*}$ the true model parameters and $x_{t}$ is the arm selected at round $t$ by the bandit algorithm employed to make decisions. This definition of regret translates directly into wasted profit and fertilizer input, making it a practical metric for resource-efficient agricultural decision-making. For each run, we record cumulative regret, average profit, and select fertilizer rates. In order to quantify the uncertainty, we also plot the distributions of cumulative regret and average profit over 10 replications at round 10, 20, and 30, respectively.
Results and Interpretation.
Figure 1 presents the average cumulative regret and the average profit per round for each algorithm. Figure 2 presents boxplots of the cumulative regret and average profit for each algorithm, evaluated at rounds 10, 20, and 30, respectively, across 10 independent replications. Several trends are immediately apparent:
  • Superiority of nonlinear model-based methods: Both nonparametric (kNN-UCB) and linear (LinUCB) methods incur substantially higher regret than the nonlinear model-based approaches. In particular, kNN-UCB and LinUCB accumulate regret rapidly and fail to converge to near-optimal recommendations within the small sample budget. LinUCB, although it does not experience a sharp initial dip in profit due to exploration, tends to settle on a suboptimal arm and consistently incurs losses relative to the optimal strategy.
  • Exploration versus exploitation trade-off: All algorithms that actively explore—notably $\epsilon$-greedy and UCB—show a pronounced dip in profit at early rounds, reflecting the cost of exploratory actions. In contrast, the ViOlin algorithm, which is designed to be more exploitative (greedily maximizing predicted profit with curvature matching), avoids this initial dip and achieves high profits almost immediately. However, both $\epsilon$-greedy and UCB eventually recover and converge to policies with low cumulative regret, validating the effectiveness of their exploration in learning the optimal fertilizer rate.
  • Uncertainty Quantification: Figure 2 highlights important differences between model-based approaches (e.g., ViOlin, model based UCB) against linear and nonparametric benchmarks such as lin-UCB and kNN-UCB, particularly in the small-sample regime relevant to agricultural field trials. While kNN-UCB exhibits relatively low variability across replicates, suggesting stable short-term performance, this stability arises primarily from smoothing rather than from capturing the underlying input-response mechanism. As a result, kNN-UCB can suffer from persistent bias, leading to consistently suboptimal mean performance. This is evident in the profit distributions, i.e., the mean profit for kNN-UCB consistently lies below the lower quartile of the nonlinear model-based approaches. In other words, the algorithm produces profits that are reliably clustered, but around a lower and biased center. This suggests that while its variability is small, its systematic bias leads to consistently suboptimal performance. In contrast, nonlinear model-based algorithms exhibit higher run-to-run variability due to parameter estimation uncertainty, particularly in early rounds when data are scarce. However, these algorithms leverage structural assumptions that align with biological processes, enabling faster convergence to the true optimum as more data accrue.
Similar trends were observed across other values of the fertilizer price $p_{x}$ and for all nonlinear response models considered. Taken together, these simulation studies demonstrate a consistent advantage of nonlinear model-based bandit algorithms over both nonparametric and linear parametric alternatives in small-sample regimes. Based on these findings, we subsequently focus our comparisons on the family of nonlinear model-based algorithms, investigating their relative strengths under a range of simulation scenarios to clarify which approaches are most advantageous in which settings.
Figure 1: Average cumulative regret and profit trajectories across 10 replications under the quadratic-plateau model ($p_{x}=0.7$). Nonlinear model-based algorithms consistently outperform linear and nonparametric baselines.
Figure 2: Uncertainty quantification in different Bandit algorithms for profit maximization/regret minimization showing how model-based methods more quickly converge to cost-efficient fertilizer recommendations.
(a) Refer to caption
(b)Refer to caption
Figure 3: Selected fertilizer rates over time for the three non-linear model based bandit algorithms
In this next part of the simulation study, we focus on illustrating the implications of using the three non-linear model-based bandit algorithms for learning fertilizer rates sequentially with the goal of profit maximization. Again in a small-sample regime, we run each of our algorithms for $T=30$ rounds and replicate each run 10 times. Since our algorithms choose arms over time, it is important to visualize how the arm choices evolve over time for each of the algorithms. In Figure 3(a), we plot the running proportion of fertilizer rate ranging from $\{0,50,\ldots,250\}$ lb N/ac, selected at each round by each of the algorithm over the 10 replicates for the quadratic plateau model for $p_{x}=\$0.5$/lb N to mimic realistic prices per unit nitrogen fertilizer (Urea) in the Midwest US. Note that in the beginning, the UCB and $\epsilon$-greedy ($\epsilon_{t}=t^{-1.5}$) explore other arms such as $x=100,150$ lb N/ac, but just after a few rounds of exploration, learn the better arm to be $x=200$ lb N/ac. ViOlin is more greedy from the beginning itself as can be seen by the proportions of other arms (other than purple) selected in the initial rounds. In Figure 3(b), in order to assess the affect of increasing fertilizer prices on decision-making for profit maximization, we also plotted the average selected fertilizer rate over time for the three algorithms. We note that the optimum arm (fertilizer rate) choice decreases as $p_{x}$ (price per unit) increases from \$0.3 to \$0.7/lb N. Additionally, since our decisions depend significantly on how well our parameters for the non-linear model are estimated over time, in Figure 4, we also plot the parameter trajectories over time for the quadratic plateau model with $p_{x}=\$0.5$/lb N. Note that all the three algorithms, as data accumulates, the parameter estimates stabilize close to the true parameter values that were used to generate the data. In terms of interpretability, one can, for example, look at the parameter estimates for $x_{0}$ (bottom right), which denotes the threshold value at which the quadratic model plateaus, thus reflecting the saturating point over rounds (or seasons), beyond which additional fertilizer amount provide little benefit. In our simulations, we also examined the sensitivity of parameter estimation to initialization. Because nonlinear least-squares fitting is iterative, the choice of starting values can influence convergence, particularly in small-sample regimes. To reflect realistic practice, we initialized parameters at values close to but not equal to the true generating parameters (see Section D in the Appendix for explicit values used in each model). We observed that biologically plausible initializations led to faster and more stable convergence, while poor starting values could slow parameter recovery. In practice, we recommend using domain knowledge (e.g., agronomic ranges for maximum yield, response rates, or plateau points) to set initial values when fitting nonlinear models. We conducted similar experiments for the other non-linear models and those results are presented in the Appendix.
Figure 4: Parameter trajectories for the quadratic plateau model over rounds
5.3 Misspecified setting
Simulation Setup.
To mimic a realistic situation in which the decision-maker fits an approximate but shape-compatible model, we generated data from the Mitscherlich response
$Y_{\mathrm{true}}(x)=A\!\left(1-e^{-b(x-d)}\right),\qquad A=120,\;b=0.015,\;d=80,$
yet fitted a quadratic-plateau curve
$Y_{\mathrm{fit}}(x)=a+bx+cx^{2}\;\text{for }x\leq x_{0},$
truncated at $x_{0}=180$ and takes the same value $a+bx_{0}+cx_{0}^{2}$ for $x>x_{0}$. All other ingredients are identical to the well-specified experiment except that we increase the time horizon to $T=100$, specifically, price grid $p_{x}\in\{0.3,0.5,0.7\}$ \$ lb-1 N, yield price $p_{y}=\$5.00$ bu-1, action set $\mathcal{X}=\{0,50,100,\ldots,250\}$ lb N ac-1, noise $\sigma=0.5$, and the 5 algorithms: $\epsilon$-greedy, nonlinear-UCB, ViOlin, LinUCB, kNN-UCB. 10 independent replicates were run for each algorithm and price.
Figure 5: Average cumulative regret trajectories across 10 replications in the misspecified setting. Nonlinear model-based algorithms still consistently outperform linear and nonparametric baselines.
(a) Refer to caption
(b)Refer to caption
Figure 6: Selected fertilizer rates over time for the three non-linear model based bandit algorithms for the misspecified setting
Figure 5 plots the average cumulative regret when the fitted quadratic-plateau and Michaelis Menten model is misspecified with respect to the true Mitscherlich process. Relative to the well-specified case, the model-based regret curves shift upwards, reflecting the price of fitting an approximate response surface. Yet their slopes remain moderate, indicating that the fitted quadratic-plateau still guides the search toward profitable regions quickly. Note, that similar to the well-specified setting, the model-based non-linear bandits perform better than the linear and non-parametric baselines. Although the quadratic-plateau form cannot capture the exponential approach to an asymptote exactly, its concave-then-flat shape matches the broad geometry of the Mitscherlich curve. As a consequence, all three nonlinear model-based algorithms, $\epsilon$-greedy, UCB, and ViOlin continue to accumulate substantially less regret than the linear (LinUCB) or non-parametric (kNN-UCB) baselines. These results highlight the resilience of shape-compatible models under misspecification and point to rich opportunities for deeper theoretical and empirical study. Although ViOlin’s greedier, curvature-penalized score again avoids the large initial dip seen in UCB and $\epsilon$-greedy, note that UCB and $\epsilon$-greedy seem to perform comparably, especially in the Michaelis Menten fit, perhaps suggesting that exploration might help in misspecified settings. Other details can be found in the Appendix.
Figure 7: Parameter estimate trajectories for the misspecified setting
Similar to the well-specified setting, we also plot the running proportion of arms over time and the most frequently selected arm trajectories in Figure 6 (a) and (b) respectively. Although the arm selections remain mostly similar as in the well-specified settings, we do notice that UCB tends to explore more in the misspecified setting (with the same exploration parameter $\alpha=1$ as in the well-specified setting), which also results in better parameter estimates for UCB as is seen in Figure 7.
These findings underline a pragmatic principle for on-farm experimentation: when the shape of the assumed response is qualitatively correct, e.g. monotone-increasing with a plateau, model-based bandits remain sample efficient even if the functional form is wrong. Conversely, generic linear or purely non-parametric methods may demand far more data before converging, a luxury seldom available to smallholders. Hence, coupling modest domain knowledge with sequential decision-making promises a robust path toward Sustainable Development Goal 2 by enabling data-scarce producers to improve input efficiency without costly large-scale trials.
6 Conclusion and discussion
In this work, we developed and systematically evaluated model-based nonlinear bandit algorithms for optimizing input decisions in agricultural systems. By casting the nitrogen rate optimization task as a sequential decision problem, we demonstrated that mechanistic model-based approaches, such as those relying on Mitscherlich, quadratic plateau, Michaelis-Menten, or logistic dose-response curves, can be embedded into adaptive learning algorithms. This integration yields decision rules that are not only sample-efficient but also interpretable, biologically grounded, and directly actionable for practice.
Our simulation studies revealed two central insights. First, under well-specified conditions, model-based bandits achieve substantially lower regret than linear or nonparametric benchmarks, underscoring the advantage of exploiting mechanistic structure. Second, under model misspecification, regret inevitably increases, but mechanistic models still provide competitive performance when the assumed shape is qualitatively correct, outperforming generic baselines in small-sample regimes. This robustness is critical in agriculture, where each additional trial represents both a financial cost and an environmental burden.
A key theoretical contribution of this work is the establishment of dimension-independent sample complexity guarantees for the ViOlin algorithm. Because each field trial corresponds to real costs of time, labor, and inputs, the ability to guarantee learning under strict resource constraints is as valuable as the specific rate of convergence. We show that the number of rounds required to identify a near-optimal fertilizer rate scales polynomially with the desired accuracy and is independent of the size of the action space. These guarantees provide a principled baseline for nonlinear model-based bandits, ensuring that learning remains sample-efficient even in high-dimensional decision settings. These bounds illustrate the theoretical tractability of model-based approaches for structured agronomic models and establish a foundation for sharper, model-specific analyses. Importantly, our simulations suggest that much faster convergence is possible in practice, underscoring a fruitful gap between theory and application that motivates further research.
Our findings highlight model choice and robustness as equally critical to algorithmic performance. The trade-off between interpretability and flexibility is especially acute in real-world sequential decision-making: mechanistic models provide transparency and scientific grounding, while more flexible models offer resilience to misspecification. We emphasize that collaboration with domain experts is indispensable for resolving this tension. Agronomic knowledge can inform priors, parameter ranges, and plausible initial values, while also guiding the interpretation of fitted parameters. Such integration ensures that adaptive management strategies are not only statistically principled but also credible and actionable for practitioners.
Looking forward, several avenues offer promising directions for future research. While our results are based on controlled simulations, future work will involve applying these methods to field experiments and real-time on-farm decision making. Such studies are essential for validating the algorithms under practical constraints, including delayed feedback, heterogeneity across fields, and limited opportunities for intervention. Developing principled strategies for model selection or model averaging within the bandit setting could mitigate misspecification risks. Extending the framework to multi-dimensional or spatially varying treatments (e.g., variable-rate nitrogen across heterogeneous fields) would better reflect on-farm conditions. Incorporating weather, soil, and other covariates into nonlinear reward models, and designing algorithms that adapt to non-stationary environments, are further priorities. Methodologically, bridging mechanistic and machine-learning approaches through hybrid or semi-parametric models could combine interpretability with flexibility. From a theoretical perspective, refining regret bounds for structured nonlinear models and extending them to settings with delayed or censored rewards would strengthen the practical utility of these algorithms.
From a broader perspective, this work contributes to the development of sample-efficient, interpretable, and robust sequential decision-making methods that can directly inform precision agriculture. By enabling smallholder farmers and resource-constrained producers to learn from limited experiments, such methods support progress toward Sustainable Development Goals, in particular SDG 2 (Zero Hunger) and SDG 12 (Responsible Consumption and Production). Overall, our study lays the groundwork for adaptive experimentation strategies that advance both the methodological frontiers of sequential decision-making while contributing to the societal goals of sustainable, data-efficient agriculture.
Reproducibility statement
All simulations were implemented in Python 3.11 using standard scientific libraries (NumPy, SciPy, Matplotlib). We report the data-generating processes, parameter values, and hyperparameters in full within the paper and Appendix, including initial values used for nonlinear least-squares estimation, fertilizer and grain price settings, and noise levels. Each experiment was repeated independently across 10 replicates with fixed randomization schemes to assess variability. Our simulation code (anonymized for review) will be made publicly available upon publication to facilitate exact replication of results and adaptation to new scenarios.
Data availability statement
The simulation code and data used in this study will be made publicly available on Github upon publication.
Ethics statement
This study does not involve human participants, human data, or animals, and therefore ethics approval was not required.
Funding statement
This research received no specific grant from any funding agency in the public, commercial, or not-for-profit sectors.
Conflict of interest
The authors declare that there are no conflicts of interest regarding the publication of this paper.
Appendix
Appendix A Rademacher complexity for parametric nonlinear function classes
Proof of Theorem 2.
We follow the approach of Rakhlin et al., 2015b . Let $N_{1}(\alpha,\mathcal{F},T)$ denote the sequential covering number in the $\ell_{1}$-norm at scale $\alpha$. From equation (9) of their paper, we have:
$\mathfrak{R}_{T}(\mathcal{F})\leq\alpha+\sqrt{\frac{2\log N_{1}(\alpha,\mathcal{F},T)}{T}}.$ (9)
Using Corollary 1 from the same paper, the $\ell_{\infty}$-sequential covering number satisfies:
$N_{\infty}(\alpha,\mathcal{F},T)\leq\left(\frac{2eT}{\alpha}\right)^{\mathrm{fat}_{\alpha}(\mathcal{F})},$
and since $N_{1}(\alpha,\mathcal{F},T)\leq N_{\infty}(\alpha,\mathcal{F},T)$, we obtain:
$\log N_{1}(\alpha,\mathcal{F},T)\leq\mathrm{fat}_{\alpha}(\mathcal{F})\cdot\log\left(\frac{2eT}{\alpha}\right).$ (10)
For any class $\mathcal{F}\subseteq[-B,B]^{\mathcal{X}}$, we have from Theorem 10 of Rakhlin et al., 2015a :
$\mathrm{fat}_{\alpha}(\mathcal{F})\leq\left\lceil\frac{2B}{\alpha}\right\rceil.$ (11)
Now, using the bounds in (10) and (11) in (9) and choosing $\alpha=\frac{B}{\sqrt{T}}$, we get:
$\left\lceil\frac{2B}{\alpha}\right\rceil=\left\lceil 2\sqrt{T}\right\rceil\leq 3\sqrt{T},\quad\log\left(\frac{2eT}{\alpha}\right)=\log\left(\frac{2eT^{3/2}}{B}\right)=\mathcal{O}(\log T).$
Thus,
$\mathfrak{R}_{T}(\mathcal{F})\leq\frac{B}{\sqrt{T}}+\sqrt{\frac{6\sqrt{T}\cdot\log T}{T}}=\frac{B}{\sqrt{T}}+\sqrt{\frac{6\log T}{\sqrt{T}}}.$
The dominant term is $\frac{B\sqrt{\log T}}{\sqrt{T}}$, so we conclude:
$\mathfrak{R}_{T}(\mathcal{F})\leq C\cdot B\cdot\sqrt{\frac{\log T}{T}}$
for some universal constant $C$. ∎
Appendix B Finding optimal profit maximizing arms for different non-linear models
Recall that $\Pi(x)$ expresses the trade-off between revenue from yield and the cost of input. To find the economically optimal fertilizer dose $x^{*}$, we differentiate the profit function with respect to $x$: Below, we derive $x^{*}$ for each model considered in this work.
Mitscherlich Model:
The Mitscherlich yield response is
$Y(x)=A(1-e^{-bx}),$
so the profit function is
$\Pi(x)=p_{y}\,A(1-e^{-bx})-p_{x}x.$
The optimal dose $x^{*}$ solves
$\frac{d\Pi}{dx}=p_{y}\,A\,b\,e^{-bx}-p_{x}=0,$
giving
$x^{*}=-\frac{1}{b}\ln\left(\frac{p_{x}}{p_{y}Ab}\right).$
Note: If $\frac{p_{x}}{p_{y}Ab}\geq 1$, then $x^{*}=0$.
Quadratic Model with Threshold
Suppose
$Y(x)=\begin{cases}a+bx+cx^{2},&x\leq x_{0}\\ a+bx_{0}+cx_{0}^{2},&x>x_{0}\end{cases}$
with threshold $x_{0}$. Then
$\Pi(x)=p_{y}Y(x)-p_{x}x.$
For $x\leq x_{0}$, set the derivative to zero:
$\frac{d\Pi}{dx}=p_{y}(b+2cx)-p_{x}=0\implies x^{*}=\frac{1}{2c}\left(\frac{p_{x}}{p_{y}}-b\right).$
Thus,
$x^{*}=\min\left\{x_{0},\,\max\left\{0,\,\frac{1}{2c}\left(\frac{p_{x}}{p_{y}}-b\right)\right\}\right\}.$
If $x>x_{0}$, profit decreases with increasing $x$.
Michaelis-Menten (MM) Model
With
$Y(x)=\frac{ax}{b+x},$
the profit is
$\Pi(x)=p_{y}\,\frac{ax}{b+x}-p_{x}x.$
Setting the derivative to zero,
$\frac{d\Pi}{dx}=p_{y}\,\frac{ab}{(b+x)^{2}}-p_{x}=0\implies(b+x)^{2}=\frac{abp_{y}}{p_{x}}$
$x^{*}=\sqrt{\frac{abp_{y}}{p_{x}}}-b.$
We set $x^{*}=0$ if the right side is negative.
Logistic Model
For the logistic response
$Y(x)=\frac{A}{1+e^{-B(x-C)}},$
profit is
$\Pi(x)=p_{y}\,\frac{A}{1+e^{-B(x-C)}}-p_{x}x.$
Let $u=e^{-B(x-C)}$, so $x=C-\frac{1}{B}\ln u$. Setting the derivative to zero,
$p_{y}\,A\,\frac{Bu}{(1+u)^{2}}=p_{x},$
which yields a quadratic in $u$:
$u^{2}+\left(2-\frac{Bp_{y}A}{p_{x}}\right)u+1=0.$
Let
$\gamma=\frac{Bp_{y}A}{p_{x}}.$
The positive root is
$u^{*}=\frac{\gamma-2-\sqrt{(\gamma-2)^{2}-4}}{2},$
and the optimal dose is
$x^{*}=C-\frac{1}{B}\ln u^{*}.$
If $u^{*}\leq 0$ or $x^{*}<0$, set $x^{*}=0$.
Appendix C Baseline Algorithm descriptions
We compare the non-linear model based algorithms to the nonparametric and linear baselines. Specifically, we compare these with the LinUCB and $k$-Nearest Neighbor UCB algorithm. For the sake of completeness, we describe these algorithms below.
The LinUCB algorithm (Algorithm 7) is a classic model-based bandit method that assumes a linear relationship between the reward and the action (or its features). At each round, LinUCB fits a linear regression model to the observed data (line 2) and predicts the expected yield for each fertilizer rate. It computes an upper confidence bound (UCB) for each action by adding a model-based uncertainty term, proportional to the standard error of the prediction, to the profit estimate (lines 4–5). The fertilizer rate with the highest UCB is selected for the next trial (line 6), balancing exploration of uncertain arms and exploitation of high-yield arms. While LinUCB is efficient and easy to implement, its performance can degrade if the true reward function is nonlinear or misspecified.
The kNN-UCB algorithm (Algorithm 8) is a nonparametric bandit method that does not assume a specific model for the yield response. Instead, for each candidate fertilizer rate, it estimates the expected yield by averaging the observed yields of the $k$ most similar (nearest) rates previously tried (line 4). The uncertainty in this prediction is quantified by the sample variance among these $k$ neighbors (line 5). As with LinUCB, an upper confidence bound is constructed for each candidate rate (line 6), and the rate with the highest UCB is selected (line 7). Although kNN-UCB is robust to model misspecification and can capture complex nonlinearities, it generally requires more data to achieve accurate estimates, particularly when the action space is large or sparsely explored.
C.1 Toy Illustration of Algorithm Behavior
To make the exploration–exploitation logic more tangible, we provide a simple illustration using a Mitscherlich yield function
$Y(x)\;=\;80+120\,\big{(}1-e^{-0.015x}\big{)},$
with crop price $p_{y}=5$ per unit yield (bu-1), fertilizer cost $p_{x}=0.7$ per unit (lb-1 N), and action grid $\{0,50,100,150,200\}$ lb N ac-1. Table 3 shows four rounds of decision making under three algorithms: $\epsilon$-Greedy, UCB, and ViOlin.
Table 3: Toy demonstration of algorithm behavior (profit in \$).
Round $\varepsilon$-Greedy UCB ViOlin
Action Profit Action Profit Action Profit
1 Explore 50 541 Explore 100 610 Explore 50 540
2 Explore 150 650 Explore 150 648 Exploit 150 649
3 Exploit 150 648 Exploit 150 647 Exploit 150 648
4 Explore 50 541 UCB $\to$ 200 645 Exploit 150 647
This toy comparison highlights the distinct decision logics:
  • $\epsilon$-Greedy alternates between random exploration and exploiting the currently estimated best dose.
  • UCB incorporates an explicit uncertainty bonus, occasionally sampling higher but uncertain doses (round 4).
  • ViOlin emphasizes curvature and stability, quickly converging to the plateau dose and avoiding unnecessary exploration.
Even in this simple 4-round setting, the three strategies exhibit clearly different exploration styles. Such illustrations help convey to applied readers that algorithm choice affects not only long-run performance, but also the sequence of recommendations farmers may observe in practice.
Appendix D Additional Simulation Results
In this section, we present the simulation results for other non-linear models that we considered in the simulation experiment of Section 5. First for the well-specified setting, we present the results for the Logistic and Mitscherlich model that was fit instead of quadratic plateau model. For reproducibility, Table 4 reports the true parameter values used to generate data in our simulations, alongside the initial values supplied to the nonlinear least-squares estimation routines. These initialization choices were selected to be reasonably close to the true parameters, reflecting what might be obtained from domain knowledge or prior agronomic experiments. As discussed in Section 5, good initialization improves convergence in small-sample regimes, though all algorithms eventually stabilize near the true values as data accumulate.
Table 4: True and initial parameter values used in simulation studies.
Model True Parameters Initial Values
Mitscherlich A=120,b=0.015,d=80A=120,\ b=0.015,\ d=80 A=100,b=0.01,d=75A=100,\ b=0.01,\ d=75
Quadratic Plateau $a=80,\ b=1.2,\ c=-0.003,\ x_{0}=180$ $a=75,\ b=1.0,\ c=-0.002,\ x_{0}=160$
Michaelis–Menten a=150,b=100,d=60a=150,\ b=100,\ d=60 a=120,b=80,d=50a=120,\ b=80,\ d=50
Logistic A=120,B=0.05,C=125,d=70A=120,\ B=0.05,\ C=125,\ d=70 A=100,B=0.03,C=100,d=65A=100,\ B=0.03,\ C=100,\ d=65
Well-specified setting: Additional results
The same hyperparameter choices were made for the algorithms as in the well-specified setting in Section 5. Namely, we use $\epsilon_{t}=t^{-1.5}$ to encourage exploitation sooner than later, $\alpha=1$ for UCB and $\alpha_{1}=2,\alpha_{2}=640$ for ViOlin. Each algorithm was run 10 times and $T=30$ to emulate data-limited setting.
Figure 8: Cumulative Regret curves for the Logistic and Mitscherlich model.
(a) Refer to caption
(b)Refer to caption
Figure 9: Selected fertilizer rates over time for the three non-linear model based bandit algorithms for the well-specified setting corresponding to Logistic and Mitscherlich models, respectively. Note that the choice of arms changes based on the underlying true model.
It can be noted from Figure 9 that while the arm choices remain mostly similar in the Logistic model to that of Quadratic plateau model, for the Mitscherlich model the ViOlin algorithm seems to choose $x=250$ ln N/ac more than other choices, which seems also reflect in the worsened regret for ViOlin in Figure 8. We hypothesize that this is because, as the yield curve quickly saturates, the local curvature used for exploration in ViOlin becomes very small, causing the algorithm to underestimate uncertainty and prematurely stop exploring. Also, we notice greater variability in the estimation of parameters especially in the Logistic model as can be seen in Figure 10.
Figure 10: Parameter estimate trajectories for the well-specified setting for the Logistic model
(a) Refer to caption
(b)Refer to caption
Figure 11: Selected fertilizer rates over time for the three non-linear model based bandit algorithms for the misspecified setting. Note that the choice of fertilizer rate decreases for both UCB and ViOlin as the price per unit fertilizer increases from 0.3 to 0.7\$ lb-1N.
Misspecified setting: Additional results
For the misspecified setting, for the same combination with the true model being Mitscherlich and fitted model being Quadratic plateau, we illustrate the effect of changing the level of fertilizer price from \$0.3 to \$0.7 per unit lb N. The same hyperparameter choices are made as in the misspecified setting of Section 5.
Figure 12: Cumulative regret curves for the misspecified model corresponding to $p_{X}=0.3$ and $0.7$ respectively.
Figure 11 (a) corresponds to $p_{X}=0.3$ while Figure 11 (b) corresponds to $p_{X}=0.7$. Note that, both UCB and ViOlin capture that as the price increases the most frequently chosen arm over time becomes $x=150$ (brown dotted line) instead of $x=200$ (purple dashed dotted line). This sort of adaptive learning is a robustness check on how effective the decision-making is towards the goal of profit maximization. However, $\epsilon$-greedy failed to adapt to this change, perhaps due to a fast decaying $\epsilon_{t}$ which made it get stuck on a sub-optimal arm in the initial rounds. This is also reflected in the poor regret performance by $\epsilon$-greedy in Figure 12.